### v1

# semicircular, free Poisson and R-transform in high order

Abstract
This work builds on our previous developments regarding a notion of freeness for tensors. We aim to establish a tensorial free convolution for compactly supported measures. First, we define higher-order analogues of the semicircular (or Wigner) law and the free Poisson (or Marčenko-Pastur) law, giving their moments and free cumulants. We prove the convergence of a Wishart-type tensor to the free Poisson law and recall the convergence of a Wigner tensor to the semicircular law. We also present a free Central Limit Theorem in this context. Next, we introduce a tensorial free convolution, define an $R$-transform, and provide the first examples of free convolution of measures.
1 Introduction
This work follows [12] and we refer to that previous study for a more detailed introduction to freeness for tensors. In this paper, we further develop and expand upon this notion of tensorial freeness.
The foundational and profoundly deep notion of freeness was introduced several decades ago by Dan-Virgil Voiculescu [33, 35]. This groundbreaking concept has given rise to a substantial body of work, notably through the contributions of Speicher, Nica, and Mingo, who have significantly advanced the development of the theory [28, 27, 26]. Freeness has found numerous applications across diverse fields, including (quantum) group theory, quantum information, statistical inference and many others, see for instance [18, 34, 16, 6].
One of the most powerful tools provided by freeness is free convolution, which describes the distribution of the sum (or product) of two freely independent elements in terms of the individual distributions of each element [32, 30, 9]. Our primary goal in this paper is to extend the concept of free convolution to the setting of tensorial freeness. From there, we can explore its implications for specific distributions. In particular, two landmark distributions in free probability theory are the semicircular (or Wigner) law and the free Poisson (or Marčenko-Pastur) law. Freeness plays a crucial role in the study of random matrices, a key area within the ecosystem of noncommutative probability spaces [27, 1, 3]. Semicircular and free Poisson laws arise notably as limiting distributions for large Wigner and Wishart matrices respectively, see [36, 25, 29, 5, 13].
In many respects, random tensors generalize random matrices to higher orders. This field is currently a vibrant area of research, driven by numerous applications in physics and computer science [2, 23, 10, 4, 22]. Hence, in this paper, we aim to establish analogous results in the framework of tensorial freeness. The principal contributions of this work are the following :
    We define the moments and the free cumulants associated to a tensor.
    We define the semicircular and the free Poisson laws of higher order, giving their moments and free cumulants.
    We prove the convergence of the measure of a Wishart tensor towards the free Poisson law of high order and we recall the convergence of a Wigner’s one towards the semicircular of high order and the free Central Limit Theorem.
    We develop the tensorial free additive convolution for compactly supported measures.
We do not use the term ”higher-order freeness,” as this designation is already attributed to a different concept [17]. In the rest of this introduction, we first recall briefly the background about trace invariants and non-crossing poset (formal definitions are postponed to Section 2) and then we will state our main results.
Maps of tensors.
This work will treat with real tensors that is an element of the vector space $\mathcal{E}^{N}_{p}=\mathbb{R}^{N}\otimes\cdots\otimes\mathbb{R}^{N}$, where $p\geq 1$ is called the order, the $p$ different copies of $\mathbb{R}^{N}$ are called the legs and $N$ is called the dimension of the legs of the tensor. The symmetric group $S_{p}$ acts on the tensors by permutation of the legs, that is $T^{\sigma}$ is the order $p$ tensor with entries $T^{\sigma}_{i_{1},\ldots,i_{p}}:=T_{i_{\sigma(1)},\ldots,i_{\sigma(p)}},$ and a real tensor is said symmetric if $T=T^{\sigma}$ for all $\sigma\in S_{p}$. A tensor $T$ of order $p$ is represented by a map with a single vertex with $p$ boundary edges, see Figure 1. A contraction of two tensors $T_{1}$ (of order $p$) and $T_{2}$ (of order $q$) along $r$ legs of dimension $N$, defined as
$(T_{1}\star_{r}T_{2})_{i_{1},\ldots,i_{p+q-2r}}:=\sum_{1\leq j_{1},\ldots,j_{r}\leq N}(T_{1})_{i_{1},\ldots,i_{p-r},j_{1},\ldots,j_{r}}(T_{2})_{j_{1},\ldots,j_{r},i_{p-r+1},\ldots,i_{p+q-2r}},$
is then a map with two vertices obtained by forming edges between the contracted legs of the tensors. We can contract along the legs we want by acting with a $\sigma$ on $T$. We used here the notation of [7]. More generally, let $\mathfrak{b}$ be a map with vertex set $V$, edge set $E$ and with $q\geq 1$ boundary edges say $\partial=(e_{1},\ldots,e_{q})$. Then if $(T_{v})_{v\in V}$ is a collection of tensors where the order of $T_{v}$ is the degree of $v$ in $\mathfrak{b}$, we can define the tensor in $\mathcal{E}^{N}_{q}$, for $i\in{{\left[\left[N\right]\right]}}^{\partial}$,
$\mathfrak{b}((T_{v})_{v\in V})_{i_{\partial}}=\sum_{i\in{{\left[\left[N\right]\right]}}^{E}}\prod_{v\in V}(T_{v})_{i_{\partial v}},$ (1)
where $\partial v$ is the sequence of neighboring edges and boundary edges of $v$. The linear combination of these maps of tensors of possibly different order encode all possible ways to contract the tensors and can be thought as the extension of the matrix polynomials in the matrix case, some examples are given in Section 2. If the map $\mathfrak{b}$ has no boundary edge we call it a trace map and for $(T_{v})_{v\in V}$ a collection of tensors where the order of $T_{v}$ is the degree of $v$ in $\mathfrak{b}$, we define the trace invariant associated to this trace map as
$\mathfrak{b}((T_{v})_{v\in V})=\frac{1}{N^{\gamma}}\sum_{i\in{{\left[\left[N\right]\right]}}^{E}}\prod_{v\in V}(T_{v})_{i_{\partial v}},$ (2)
where $\gamma$ is the number of connected components of $\mathfrak{b}$. The distribution of a collection of tensors $\mathcal{A}=\{A_{1},\ldots,A_{n}\}$ of possibly different order, $A_{k}\in\mathcal{E}^{N}_{p_{k}}$ is the collection of all trace maps $\mathfrak{b}((T_{v})_{v\in V})$ with $T_{v}\in\mathcal{A}$ and $\mathfrak{b}$ with compatible degrees. In particular, if $T\in\mathcal{E}^{N}_{p}$, the distribution of $T$ is the collection of trace invariants for $p$-regular trace maps, that is all vertices of the map belong to exactly $p$-edges.
We introduce now informally three types of $p$-regular trace maps useful for the rest of the introduction. A melon is a map with two vertices and $p$ edges between them. If $p$ is even, a bouquet is a map with one vertex and $p/2$ self-loops, and a multicycle is a map with $n\geq 1$ vertices $p/2$ edges between the $i$-th and the $i+1$-th ones for all $1\leq i\leq n$ ($n+1=1$). For $p$ odd, an odd multicycle is a map with $2n$ vertices and $(p+1)/2$ edges between the $12i-$ 1-th and the $2i$-th ones, $(p-1)/2$ edges between the $2i$-th and the $12i+$ 1-th ones, for all $1\leq i\leq n$ ($12n+1=$ 1).
Non-crossing poset.
The notion of freeness detailed in [12] arises from a poset (partially ordered set) on maps, called by analogy the non-crossing poset. Loosely speaking, we say that a map $\mathfrak{b}^{\prime}$ is smaller than a map $\mathfrak{b}$ if they differ by a switch, that is exchanging the ending point of two edges in $\mathfrak{b}$, and $\mathfrak{b}^{\prime}$ has one more connected component. By transitivity, this gives a poset on the set of trace maps with a given number of vertices and a given sequence of degrees. The formal definitions are recall in Section 2 and an example is given in Figure 2. By Moebius inversion, it is possible to define free cumulants associated to this poset characterized by the implicit relation
$\mathfrak{b}((T_{v})_{v})=\sum_{\mathfrak{b}^{\prime}\leq\mathfrak{b}}\kappa_{\mathfrak{b}^{\prime}}((T_{v})_{v}).$ (3)
There is a notion of freeness associated to this poset where importantly the distribution of a family of freely independent families is characterized by the individual distribution of the families. We also proved that the vanishing of mixed cumulants characterized freeness, but only in the case of even families for technical issues, as we will discuss later. We are now ready to go into the main results of this work.
1.1 Main results
For a tensor $T$ of order $p$, and $n\geq 1$, we define $m_{n}(T)$, the $n$-th moment of $T$, as the sum over all the $p$-regular connected rooted trace maps with $n$ vertices of the trace invariants where all vertices are decorated by $T$. Similarly, $\kappa_{n}(T)$ the $n$-th free cumulant of $T$ is the sum over all the $p$-regular connected trace maps with $n$ vertices of free cumulant of the trace map where all vertices are decorated by $T$. We set $m_{0}(T)=\kappa_{0}(T)=1$. We then define the corresponding formal power series in $\mathbb{C}{{\left[\left[z\right]\right]}}$ :
$M_{T}(z):=\sum_{n\geq 0}m_{n}(T)z^{n}\quad\quad\mathrm{and}\quad\quad C_{T}(z):=\sum_{n\geq 0}\kappa_{n}(T)z^{n}.$
Theorem 1 (Analytic moment-cumulant formula).
We have as functionnal relation in $\mathbb{C}{{\left[\left[z\right]\right]}}$,
$M_{T}(z)=C_{T}\left(zM_{T}(z)^{p/2}\right).$
Note that this relation is always in $\mathbb{C}{{\left[\left[z\right]\right]}}$ and not in $\mathbb{C}{{\left[\left[\sqrt{z}\right]\right]}}$ since if $p$ is odd, then all the odd moments are zero as there is no $p$-regular trace map with $12n+$ 1 vertices, as it would need to have $p(2n+1)/2$ edges. Moreover, Gurau showed in [21] that if $T$ is a real symmetric tensor then there exists a probability measure $\mu_{T}$ such that for all $n\geq 0$,
[…]
We have not yet succeeded to give a direct argument of the existence of this measure by showing the positivity of these moments, nevertheless the following results of convergence can always be given in terms of convergence of the moment and vanishing of the variance instead of in terms of weak convergence of the measure.
If $a$ is a distribution on the $p$-regular combinatorial maps (that is $a:\mathcal{M}_{0}\mapsto\mathbb{R}$ with $a(\emptyset)=1$), we can speak of its moments $m_{n}(a)$ as the sum of $a(\mathfrak{b})$ over all the $p$-regular connected maps with $n$ vertices, and we denote $\mu_{a}$ as the measure on $\mathbb{R}$ having $(m_{n}(a))_{n\geq 0}$ as moments, if it exists.
High order laws.
For $p\geq 1$, we define the two following distributions $a_{p}$ and $b_{p,t}$ on the $p$-regular combinatorial maps by
$a_{p}(\kappa_{\mathfrak{b}})=\frac{1}{(p-1)!}\mbox{ if }\mathfrak{b}\mbox{ is a melon and }0\mbox{ otherwise,}$,
    if $p$ is even,
    $b_{p,t}(\kappa_{\mathfrak{b}})=\frac{t}{[(p/2)!]^{n}}\mbox{ if }\mathfrak{b}\mbox{ is a multicycle of length }n\mbox{ and }0\mbox{ otherwise,}$,
    if $p$ is odd,
    $b_{p,t}(\kappa_{\mathfrak{b}})=\frac{t}{[(p+1/2)!(p-1/2)!]^{n}}\mbox{ if }\mathfrak{b}\mbox{ is an odd multicycle of length }2n\mbox{ and }0\mbox{ otherwise.}$.
Theorem 2 (Semicircular law in high order).
For $p\geq 1$, the measure $\mu_{p}:=\mu_{a_{p}}$ is a compactly supported probability measure on $\mathbb{R}$ having for moments
$m_{n}(\mu_{p})=\frac{1}{pn/2+1}\binom{pn/2+1}{n/2}\mbox{ if $n$ even}\quad\quad\mathrm{and}\quad\quad m_{n}(\mu_{p})=0\mbox{ if $n$ odd},$
and for free cumulants
$\kappa_{2}(\mu_{p})=1\mbox{ if $n=0$ or $2$}\quad\quad\mathrm{and}\quad\quad\kappa_{n}(\mu_{p})=0\mbox{ otherwise}.$
The number $F_{p}(k)=\frac{1}{pk+1}\binom{pk+1}{k}$ is the $k$-th Fuss-Catalan number of order $p$. They extend the Catalan numbers for $p\geq 2$.
Theorem 3 (Free Poisson law in high order).
For $p\geq 2$, the measure $\nu_{p,t}:=\mu_{b_{p,t}}$ is a compactly supported probability measure on $\mathbb{R}$ having for moments
    if $p$ is even,
    $m_{n}(\nu_{p,t})=\sum_{b=1}^{n}\frac{1}{b}\binom{n-1}{b-1}\binom{pn/2}{b-1}t^{b},$
    if $p$ is odd,
    $m_{n}(\nu_{p,t})=\sum_{b=1}^{n}\frac{1}{b}\binom{n-1}{b-1}\binom{pn/2}{b-1}t^{b}\mbox{ if $n$ even}\quad\quad\mathrm{and}\quad\quad m_{n}(\nu_{p,t})=0\mbox{if $n$ odd},$
and for free cumulants
    if $p$ is even,
    $\kappa_{n}(\nu_{p,t})=t\mbox{ for all $n\geq 1$},$
    if $p$ is odd,
    $\kappa_{n}(\nu_{p,t})=t\mbox{ for all even $n\geq 1$}\quad\quad\mathrm{and}\quad\quad m_{n}(\nu_{p,t})=0\mbox{ for all odd $n\geq 1$}.$
The polynomial $F_{p,k}(t)=\sum_{b=1}^{n}\frac{1}{b}\binom{k-1}{b-1}\binom{pk}{b-1}t^{b}$ is called the $k+1$-th Fuss-Narayana polynomial of order $p$ and parameter $t$. In particular, we have that $F_{p,k}(1)=F_{p-1}(k)$. Interestingly, we find relations between the high order semicircular and free Poisson of parameter $t$ when $t=1$ and when $t\rightarrow\infty$, extending the ones known for $p=2$, see later.
We may also define $\omega_{p,t}$ the Marčenko-Pastur law of parameter $\tau$ as the free Poisson law of parameter $1/\tau$ dilated by a factor $1/\tau^{p/2}$. In the matrix case, that is associated to the densities
$d\nu_{2,t}(x)=(1-t)_{+}\delta_{0}+\frac{\sqrt{4t-(x-1-t)^{2}}dx}{2\pi x}dx,$
$d\omega_{2,t}(x)=(1-\frac{1}{\tau})_{+}\delta_{0}+\frac{\sqrt{4\tau-(x-1-\tau)^{2}}}{\tau\times 2\pi x}dx,$
which have both moments given by some Narayana numbers. This is due to a symmetry in these numbers changing $b$ in $n-b$ which does not subsist when $p>2$, see Remark 3 for more details. That is why we chose this convention in the definitions of free Poisson and Marčenko-Pastur laws.
Convergence results.
The analogy between these laws in high order and the usual ones is not simply justified by the extension of the moments and free cumulants to higher $p$. We have also the main classical results of convergence which remains true. In particular, we have convergence of a Wigner tensor towards the higher order semicircular law, convergence of a Wishart tensor towards the higher order free Poisson law, and a free Central Limit Theorem.
Formal definitions of a Wigner and Wishart tensor will come in Section 3.1. Just in some words, a Wigner tensor is of the form
$\mathbf{W}_{N}:=\frac{X}{N^{\frac{p-1}{2}}},$
where $X$ is a real symmetric tensor with entries mean 00, variance $\frac{1}{(p-1)!}$ and bounded moments. This is in particular the case when the entries are Gaussian. On the other hand, a Wishart tensor is a real symmetric tensor of the form
$\mathcal{W}^{k}_{N}:=\frac{x_{1}^{\underline{\otimes}2}+\ldots+x_{k}^{\underline{\otimes}2}}{N^{\frac{p}{2}}},$
where the $x_{i}$ are $(p/2)$-order tensors with i.i.d entries with mean 00, variance $\frac{1}{[(p/2)!]^{2/p}}$ and bounded moments (it is for even $p$, see odd case and formal definitions in Section 3.1). We then have the following convergences.
Theorem 4 (High order Wigner Theorem).
Let $\mathbf{W}^{p}_{N}$ be a $p$-order Wigner tensor and $\mu_{\mathbf{W}_{N}}$ its associated measure. Then, when $N\rightarrow\infty$, we have weak convergence in probability,
$\mu_{\mathbf{W}_{N}}\rightarrow\mu_{p}.$
In other words, we are going to prove that when $N\rightarrow\infty$,
$\mathbb{E}[m_{n}(\mathbf{W}_{N})]=m_{n}(\mu_{p})+\mathcal{O}(1/N)\quad\quad\mathrm{and}\quad\quad\mathrm{Var}[m_{n}(\mathbf{W}_{N})]=\mathcal{O}(1/N^{2}).$
Theorem 5 (High order Marčenko-Pastur Theorem).
Let $\mathcal{W}^{p,k}_{N}$ be a $p$-order Wishart tensor such that $k_{N}/N^{p/2}\rightarrow t\in(0,\infty)$ and $\mu_{\mathcal{W}_{N}}$ its associated measure. Then, when $N\rightarrow\infty$, we have weak convergence in probability,
$\mu_{\mathcal{W}_{N}}\rightarrow\nu_{p,t}.$
In other words, we are going to prove that when $N\rightarrow\infty$,
$\mathbb{E}[m_{n}(\mathcal{W}_{N})]=m_{n}(\nu_{p,t})+\mathcal{O}(1/N)\quad\quad\mathrm{and}\quad\quad\mathrm{Var}[m_{n}(\mathcal{W}_{N})]=\mathcal{O}(1/N^{2}).$
Theorem 5 will be proven in Section 3.3. Theorem 4 has already been proved in [21] in the Gaussian case and in [11] for the general case. We will give a reminder of the ideas of the proof in Section 3.2. We note that a stronger result can in fact be stated, that is that the convergence is in distribution (at the level of the individual maps), which is given in [[11] Lemma 3.16] for Wigner, and can be easily deduced from the proof for Wishart. This is the content of the following Lemma 1, but will not be our interest in this paper as we consider the moments and not the distribution of a tensor.
Lemma 1.
We have convergence in distribution of $\mathbf{W}^{p}_{N}$ (respectively $\mathcal{W}^{p,k}_{N}$) towards $a_{p}$ (respectively $b_{p,t}$), that is for all trace map $\mathfrak{b}$, $\mathfrak{b}(\mathbf{W})\rightarrow a_{p}(\mathfrak{b})$ (resp. $\mathcal{W}$ and $b$).
The last convergence result we state is the free Central Limit Theorem for tensors.
Theorem 6 (Free CLT).
Let $p\geq 2$ be a fixed even integer and $(T_{i})_{i\geq 1}$ a collection of tensors of order $p$ such that
$m_{1}(T)=0\quad\quad\mathrm{and}\quad\quad m_{2}(T)=1\quad\quad\mathrm{and}\quad\quad\forall i,\forall\mathfrak{b}\in\mathcal{M}_{0},|\mathfrak{b}(T_{i})|\leq C(\mathfrak{b}).$
Denote $T^{k}=\frac{1}{\sqrt{k}}\sum_{i=1}^{k}T_{i}$. Then when $k$ goes to infinity,
$m_{n}(T^{k})\rightarrow m_{n}(\mu_{p}).$
We initially presented this result in [12] (Section 3.3), but not considering the sum over connected maps having the same number of vertices. We will provide it in this setting in Section 3.
Free convolution of tensors.
Theorem 1 allows us to compute the law of the sum of two semicircular or free Poisson freely independent. For the rest of this introduction, we fix $p\geq 2$ an even integer, as the following proofs rely on the vanishing of mixed cumulants.
Corollary 1.
The free convolution of two freely independent semicirculars of order $p$ is a semicircular of order $p$ dilated by a factor $\sqrt{2}$. That is,
$\mu_{p}\oplus_{p}\mu_{p}=\mu_{p}^{(\sqrt{2})}.$
Corollary 2.
The free convolution of two freely independent free Poisson of order $p$ and parameters $t$ and $t^{\prime}$ is a free Poisson of order $p$ and parameter $t+t^{\prime}$. That is,
$\nu_{p,t}\oplus_{p}\nu_{p,t^{\prime}}=\nu_{p,t+t^{\prime}}.$
We can also define the $R$-transform which is additive for two freely independent tensors.
Theorem 7.
The $R$-transform and the $Q$-transform are defined as
$R_{\mu}(z):=\frac{C_{\mu}(z)-1}{z}\quad\quad\mathrm{and}\quad\quad Q_{\mu}(z):=\frac{C_{\mu}(z)^{p/2}-1}{z}.$
For $\mu,\nu$ two compactly supported measures we have
$R_{\mu\oplus_{p}\nu}(z)=R_{\mu}(z)+R_{\nu}(z),$
for all $|z|$ sufficiently small.
We point out that if the measures associated to some models of real symmetric tensors converge to compactly supported probability measure, the measure associated to a given tensor has in general a support on $\mathbb{R}$, so we must develop subordination functions and other analytic tools to deal with free convolution in a general setting. It is where the $Q$-transform could be useful. These developments will be the subject of further works, jointly with a better understanding of the link between this measure and the $z$-eigenvalues of a given tensor.
These results also open paths to treat some questions about high dimension statistics for random tensors. One example of such an application of free convolution, in the matrix case, is given by the work done in [14]. The author would like to inquire about them in the future.
1.2 Organization of the paper
In Section 2, we will introduce some formalism and prove the analytic moment-cumulant formula. In Section 3, we will define Wigner and Wishart and prove the convergence of their associated laws. In Section 4, we will finally introduce tensorial free convolution for compactly supported measures, $R$-transform and give some basic examples.
1.3 Acknowledgements
The author wishes to extend his warm gratitude to his thesis advisors, Charles Bordenave, with whom he collaborates most of the time in Marseille, and Djalil Chafaï, with whom he works in Paris, for their guidance, availability, and the wealth of explanations, feedback, and ideas shared during their inspiring discussions.
2 Moments and free cumulants for tensors
An important difference is that when [12] and [24] only consider the distribution of the trace invariants individually, we will here define the moments of a tensor on a coarser collection of invariants as done in [21] and [11].
2.1 Maps of tensors
For integer $n\geq 1$, we set ${{\left[\left[n\right]\right]}}=\{1,\ldots,n\}$.
Combinatorial maps.
We recall some notions already introduced in [12] and we refer to it for more details. The formalism of combinatorial maps is a way to encode finite graphs $\mathfrak{b}$ equipped with an order of edges attached to each vertex, by a pair of permutations. For $m$ even integer, a combinatorial map $\mathfrak{b}=(\pi,\alpha)$ with $\pi,\alpha\in S_{m}$ has $m/2$ edges and has vertex set $V(\mathfrak{b})={{\left[\left[n\right]\right]}}$ where $n$ is the number of cycles of $\pi$. The set $\vec{E}(\mathfrak{b})={{\left[\left[m\right]\right]}}$ are the directed edges (or half-edges), $\pi$ has $n$ cycles ordered by least elements which are the directed edges attached to each vertex and $\alpha$ is an involution without fixed point whose $m/2$ cycles of length $2$ are identified as $E(\mathfrak{b})$ the edges of $\mathfrak{b}$. For $v\in V$, we denote by $\partial v=(e_{1},\ldots,e_{p})\in\vec{E}(\mathfrak{b})^{p}$ the cycle of $\pi$ associated to $v$. We always choose $e_{1}$ such that $e_{1}=\min\partial v$. The degree of $v\in V$, $\deg(v)$ is the length of the cycle, that is $p$. We denote by $\mathcal{M}_{0}$ the set of such combinatorial maps with no boundaries, which we call trace maps.
It is also possible to introduce $\mathcal{M}_{q}$ the set of combinatorial maps with $q$ boundaries, with the only difference that $m$ is no more necessarily even and $\alpha$ has $q$ fixed points which are the boundaries of the maps. A tensor $T$ of order $p$ is then represented has a map with one vertex decorated by $T$ and $p$ boundary edges. The set of all combinatorial maps is denoted $\mathcal{M}=\sqcup_{q}\mathcal{M}_{q}$. This will not be the center of our interest in this work so we do not give more details.
Some examples are given in Figure 1. In particular, we describe some simple maps appearing in the sequel. For $p\geq 1$, the melon maps of degree $p$ (or Frobenius pair in the terminology of [24]), denoted $\mathfrak{f}_{p}^{\sigma}$ for some $\sigma$, are the maps with two vertices and $p$ edges between them : $\pi=(1,\ldots,p)(p+1,\ldots,2p)$ and $\alpha=(1,p+\sigma(1))\cdots(p,p+\sigma(p))$. There are $(p-1)!$ such maps non equivalent, the canonical one being again $\mathfrak{f}_{p}^{\mathrm{id}}$. For $p=2t$ even, the bouquet maps, denoted $\mathfrak{b}_{p}^{\sigma}$ for some $\sigma$, are the ones with a single vertex and $t$ self-loops : $\pi=(1,\ldots,p)$ is a cycle and $\alpha=(\sigma(1),\sigma(2))\cdots(\sigma(p-1),\sigma(p))$. The canonical one is $\mathfrak{b}_{p}^{\mathrm{id}}$. It is a particular case of multicycle map, that is $n\geq 1$ vertices $p/2$ edges between the $i$-th and the $i+1$-th ones for all $1\leq i\leq n$ ($n+1=1$). For a given choice of inputs and outputs edges, there are $[(p/2)!]^{n}$ such maps. For $p$ odd, the definition of an odd multicycle, presented in the introduction follows.
$M$$\frac{1}{N}\mathrm{Tr}(M)$
[…]
$\mathfrak{b}_{p}^{\mathrm{id}}(T)$
$T$$T$$\mathfrak{f}_{p}^{\mathrm{id}}(T)$
$T$$U$$U$$U$$U$$T\cdot U^{p}$
$T$$u$$v$$w$$T.(u,v,w)$
Figure 1: Some tensor maps.
For a combinatorial map $\mathfrak{b}=(\pi,\alpha)$ we denote #$\#\mathfrak{b}$ the number of vertices of $\mathfrak{b}$ (number of cycles of $\pi$) and $\gamma(\mathfrak{b})$ the number of connected components if $\mathfrak{b}$ is a trace map, $\gamma(\mathfrak{b})=0$ otherwise. Furthermore, for a given $p$ we say that a map is $p$-regular if $\pi$ has only cycles of length $p$ (all vertices are $p$-valent).
Tensor maps.
For $N\geq 1$, we set $\mathcal{E}^{N}_{p}:=(\mathbb{R}^{N})^{\otimes p}$, that is $\mathcal{E}^{N}_{1}$ are real vectors, $\mathcal{E}^{N}_{2}$ are real matrices, and, for $p\geq 3$, $\mathcal{E}^{N}_{p}$ are real tensors with $p$ legs of dimension $N$. For $T\in\mathcal{E}^{N}_{p}$, we write $T=(T_{i})_{i\in{{\left[\left[N\right]\right]}}^{p}}\in\mathcal{E}^{N}_{p}$. The action of $\mathcal{M}$ on $\mathcal{E}^{N}=\sqcup\mathcal{E}^{N}_{p}$ is defined for $\mathfrak{b}\in\mathcal{M}_{q}$, with $\partial=(e_{1},\ldots,e_{q})\subset\vec{E}(\mathfrak{b})$ being its boundary, as
$\mathfrak{b}((T_{v})_{v\in V(\mathfrak{b})})_{i_{\partial}}=\frac{1}{N^{\gamma(\mathfrak{b})}}\sum_{i\in{{\left[\left[N\right]\right]}}^{E(\mathfrak{b})}}\prod_{v\in V(\mathfrak{b})}(T_{v})_{i_{\partial v}}.$ (4)
Note that we normalize by the number of connected components for the trace maps ($q=0$), and we do not normalize otherwise ($q\geq 1$).
For $\mathfrak{b}\in\mathcal{M}_{0}$, the application $\mathfrak{b}:(T_{v})_{v\in V}\to\mathbb{R}$ is multi-linear and importantly, it is orthogonal invariant in the following sense. If $T\in\mathcal{E}^{N}_{p}$ and $U\in\mathcal{E}_{2}^{N}$ is an orthogonal matrix, define $T\cdot U^{p}\in\mathcal{E}^{N}_{p}$ as the contraction of each leg of $T$ by $U$, that is for $j\in{{\left[\left[N\right]\right]}}^{p}$
$(T\cdot U^{p})_{j}=\sum_{i\in{{\left[\left[N\right]\right]}}^{p}}T_{i}\prod_{k=1}^{p}U_{j_{k}i_{k}}.$ (5)
In other words, we have $T\cdot U^{p}=\mathfrak{b}((T,U,\ldots,U))$ where $\mathfrak{b}$ is an elongated star map, with $T$ in the middle and $U$ on each branch, the second neighboring edge of $U$ being attached to $T$. If $M\in\mathcal{E}^{N}_{2}$, then $M\cdot U^{2}=UMU^{\intercal}$. Then, it is straightforward to check that for any orthogonal matrix $U$ and any trace maps,
$\mathfrak{b}((T_{v}\cdot U^{p_{v}})_{v\in V})=\mathfrak{b}((T_{v})_{v\in V}),$
where $p_{v}$ is the degree of degree $v$ (and the order of $T_{v}$). The trace maps $\mathfrak{b}$ form a basis of orthogonal invariant multi-linear application. They are the natural generalization of trace for matrices so we call them the trace invariants. We refer to [20, 24] for an introduction on these trace invariants (for tensors of even order, it is also possible to define maps which are unitary invariant). We call the distribution of a collection of tensors the collection of trace invariants for $\mathfrak{b}$ connected and adapted to the collection.
2.2 Non-crossing poset
Let $m$ be an even integer and $\pi\in S_{m}$. We consider $\mathcal{M}_{\pi}\subset\mathcal{M}_{0}$ the set of maps $\mathfrak{b}=(\pi,\alpha)$ for some $\alpha\in S_{m}$. We construct $\mathcal{G}_{\pi}$ the graph on $\mathcal{M}_{\pi}$ where two maps $\mathfrak{b}=(\pi,\alpha)$ and $\mathfrak{b}^{\prime}=(\pi,\alpha^{\prime})$ are connected by an edge if $\alpha\alpha^{\prime}$ is a product of transpositions. We say that $\mathfrak{b}$ and $\mathfrak{b}^{\prime}$ differ by a switch. Moreover, we say that $\mathfrak{b}^{\prime}<\mathfrak{b}$ if they are connected by an edge and $\gamma(\mathfrak{b}^{\prime})=\gamma(\mathfrak{b})+1$. We extend the relation $<$ by transitivity giving to $\mathcal{M}_{\pi}$ the structure of a partially ordered set. We call it by analogy the non-crossing poset as we retrieve the usual one when all vertices of the maps considered have valence $2$. Also, we say that $\mathfrak{b}$ is minimal if there is no $\mathfrak{b}^{\prime}<\mathfrak{b}$. When all the degrees are even, there is a unique minimal map smaller than a given map as the graph is Eulerian, but it is no more the case when odd degrees appear. That is why we were not able to prove the characterization of freeness by the free cumulants in the odd case. We may probably define the free cumulants of $\mathfrak{b}^{\prime}\leq\mathfrak{b}$ with a multiplicity being the number of minimal $\mathfrak{b}^{\prime\prime}$ such that $\mathfrak{b}^{\prime\prime}\leq\mathfrak{b}^{\prime}\leq\mathfrak{b}$.
[…]
Figure 2: Poset $\mathcal{P}_{\pi}$
There is an associated notion of freeness defined in [12]. Importantly, if two tensors $T_{1},T_{2}$ of possibly different orders are freely independent then the distribution of $(T_{1},T_{2})$ is characterized by the individual distributions of $T_{1}$ and $T_{2}$. In particular, we mention that we proved the asymptotic freeness for unitarily invariant families, and in particular Wigner tensors.
2.3 Moments of a tensor
If $T$ is tensor of order $p$ an $\mathfrak{b}$ a $p$-regular trace map, we may define the trace invariant
$\mathfrak{b}(T,\ldots,T)=\frac{1}{N^{\gamma(\mathfrak{b})}}\sum_{i\in{{\left[\left[N\right]\right]}}^{E}}\prod_{v\in V}T_{i_{\partial v}}.$ (6)
We denote $\mathfrak{b}(T)$ instead of $\mathfrak{b}(T,\cdots,T)$ for ease of notations, and the distribution of $T$ is then the collection of the $\mathfrak{b}(T)$ for all $p$-regular trace map.
We define $\mathcal{B}_{n}$ the set of connected rooted $p$-regular combinatorial maps in $\mathcal{M}_{0}$ with $n$ (unlabeled) vertices.
Definition 1 (Moments of a tensor).
For $T$ a tensor of order $p$, we call
$m_{n}(T):=\sum_{\mathfrak{b}\in\mathcal{B}_{n}}\mathfrak{b}(T)$
the $n$-th moment of $T$. By convention we set $m_{0}(T)=1$.
Finally, we define the generating function of these moments, for $z\in\mathbb{C}$,
$M_{T}(z):=\sum_{n\geq 0}m_{n}(T)z^{n},$
as a formal power series in $\mathbb{C}{{\left[\left[z\right]\right]}}$.
2.4 Free cumulants
Given our poset we can naturally define the associated free cumulants. For $\mathfrak{b}\in\mathcal{M}_{0}$, the free cumulant of $\mathfrak{b}$ is the unique multilinear application such that
$\mathfrak{b}((T_{v})_{v})=\sum_{\mathfrak{b}^{\prime}\leq\mathfrak{b}}\kappa_{\mathfrak{b}^{\prime}}((T_{v})_{v}).$
It is well defined by Moebius inversion. We refer to the precedent paper for all the properties of this application and their proofs. In particular note that the morphism property still holds, that is it is multiplicative on the connected components. The crucial point is that we proved that two even families of tensors of possibly different orders are free if and only if the mixed cumulants vanish.
As for the moments, we may then define the free cumulants of a given tensor $T\in\mathcal{E}_{p}$.
Definition 2 (Free cumulants of a tensor).
For $T$ a tensor of order $p$, we call
$\kappa_{n}(T):=\sum_{\mathfrak{b}\in\mathcal{B}_{n}}\kappa_{\mathfrak{b}}(T)$
the $n$-th free cumulant of $T$. By convention we set $\kappa_{0}(T)=1$.
The generating function of these free cumulants is now defined as the formal power series in $\mathbb{C}{{\left[\left[Z\right]\right]}}$,
$C_{T}(z):=\sum_{n\geq 0}\kappa_{n}(T)z^{n}.$
2.5 Moment-cumulant formula
We fix $p\geq 2$ and $T\in\mathcal{E}_{p}$. We are going to prove the Theorem 1 which gives a key to compute the moments of the sum of two tensors freely independent. That states that we have the functional relation in $\mathbb{C}{{\left[\left[z\right]\right]}}$,
$M_{T}(z)=C(zM_{T}(z)^{p/2}).$
This equality holds on the crucial relation (7) given in the following Lemma.
Lemma 2.
For any $n\geq 0$,
$m_{n}(T)=\sum_{s=1}^{n}\kappa_{s}(T)\sum_{i_{1},\cdots,i_{sp/2}\in\{0,\cdots,n-s\}\atop s+i_{1}+\cdots+i_{sp/2}=n}\prod_{j=1}^{sp/2}m_{i_{j}}(T).$ (7)
Proof of Lemma 2.
For a given $n\geq 1$, consider a map $\mathfrak{b}\in\mathcal{B}_{n}$ and a map $\mathfrak{b}^{\prime}$ smaller than $\mathfrak{b}$. Write $\mathfrak{b}_{0}$ the connected component of $\mathfrak{b}^{\prime}$ containing the vertex $1$ and denote $s$ its size. Hence $\mathfrak{b}_{0}\in\mathcal{B}_{s}$ and this map has $sp/2$ edges. The other connected components of $\mathfrak{b}_{0}$ are denoted $\mathfrak{b}_{1},\ldots,\mathfrak{b}_{r}$. By definition of the poset, for each $1\leq l\leq r$, there exists an edge of $\mathfrak{b}_{0}$ such that one edge of $\mathfrak{b}_{l}$ can be switched with this edge of $\mathfrak{b}_{0}$ and the new map obtained is still smaller than $\mathfrak{b}$ in the poset. This means that $\mathfrak{b}^{\prime}$ decomposes into $\mathfrak{b}_{0}\in\mathcal{B}_{s}$ and $\mathfrak{b}^{\prime}_{1},\ldots,\mathfrak{b}^{\prime}_{sp/2}$ (these maps are possibly empty and not necessarily connected). Therefore, using the morphism property of $\kappa_{\mathfrak{b}}$,
$\kappa_{\mathfrak{b}^{\prime}}(T)=\kappa_{\mathfrak{b}_{0}}(T)\kappa_{\mathfrak{b}^{\prime}_{1}}(T)\ldots\kappa_{\mathfrak{b}^{\prime}_{sp/2}}(T).$
If we note $i_{j}$ the number of vertices in $\mathfrak{b}^{\prime}_{j}$, we thus have proved that
$m_{n}(T)$ $=\sum_{\mathfrak{b}\in\mathcal{B}_{n}}\sum_{\mathfrak{b}^{\prime}\leq\mathfrak{b}}\kappa_{\mathfrak{b}^{\prime}}(T)$
$=\sum_{s=1}^{n}\sum_{\mathfrak{b}_{0}\in\mathcal{B}_{s}}\kappa_{\mathfrak{b}_{0}}(T)\sum_{i_{1},\cdots,i_{sp/2}\in\{0,\cdots,n-s\}\atop s+i_{1}+\cdots+i_{sp/2}=n}\prod_{j=1}^{sp/2}\sum_{\mathfrak{b}_{j}\in\mathcal{B}_{i_{j}}}\sum_{\mathfrak{b}^{\prime}_{j}\leq\mathfrak{b}_{j}}\kappa_{\mathfrak{b}^{\prime}_{j}}(T)$
$=\sum_{s=1}^{n}\kappa_{s}(T)\sum_{i_{1},\cdots,i_{sp/2}\in\{0,\cdots,n-s\}\atop s+i_{1}+\cdots+i_{sp/2}=n}\prod_{j=1}^{sp/2}m_{i_{j}}(T).$
The Lemma is proved. ∎
Proof of Theorem 1.
We inject (7) in the expression of $M_{T}(z)$,
$M_{T}(z)$ $=1+\sum_{n=1}^{\infty}\sum_{s=1}^{n}\kappa_{s}(T)\sum_{i_{1},\cdots,i_{sp/2}\in\{0,\cdots,n-s\}\atop s+i_{1}+\cdots+i_{sp/2}=n}\prod_{j=1}^{sp/2}m_{i_{j}}(T)z^{n}$
$=\sum_{s=0}^{\infty}\kappa_{s}(T)z^{s}\sum_{i_{1},\cdots,i_{sp/2}}\prod_{j=1}^{sp/2}(m_{i_{j}}(T)z^{i_{j}})$
$=\sum_{s=0}^{\infty}\kappa_{s}(T)z^{s}M_{T}(s)^{sp/2}.$
This gives the result. ∎
3 Higher order laws
3.1 Wigner and Wishart tensor
Symmetric tensors and symmetrization.
For a given $p\geq 1$, the symmetric group $\mathrm{S}_{p}$ acts on ${{\left[\left[N\right]\right]}}^{p}$ by permutation of indices: for $i\in{{\left[\left[N\right]\right]}}^{p}$ and $\sigma\in\mathrm{S}_{p}$, $i_{\sigma}=(i_{\sigma(1)},\ldots,i_{\sigma(p)})$. For $i,j\in{{\left[\left[N\right]\right]}}^{p}$, we say $i\stackrel{{\scriptstyle p}}{{\sim}}j$ if $i=j_{\sigma}$ for some $\sigma\in\mathrm{S}_{p}$. We say that a tensor is symmetric $T_{i}=T_{j}$ as soon as $i\stackrel{{\scriptstyle p}}{{\sim}}j$, that is $T^{\sigma}=T$ for all $\sigma$, where $T^{\sigma}_{i}=T_{i_{\sigma}}$.
[…]
Figure 3: $T$ and $T^{\sigma}$.
If $p_{1},p_{2}$ are two integers, we denote $\mathcal{P}_{p_{1},p_{2}}$ the set of application $\pi:{{\left[\left[p_{1}+p_{2}\right]\right]}}\mapsto\{1,2\}\times{{\left[\left[p_{1}+p_{2}\right]\right]}}$ such that $|\pi_{1}^{-1}(1)|=p_{1},|\pi_{1}^{-1}(2)|=p_{1}$ and $\pi_{2}$ is injective. We write $\pi^{-1}(1)$ the ordered $p_{1}$-tuple of elements in $\pi_{1}^{-1}(1)$ ordered by their image by $\pi$, that is $\pi^{-1}(1)=(x_{1},\ldots,x_{p_{1}})$ such that $\pi(x_{j})=(1,i_{j})$ where $i_{1}<i_{2}<\ldots<i_{p_{1}}$. We do the same for $\pi^{-1}(2)$. Now if $T_{1},T_{2}$ are tensors of order $p_{1},p_{2}$, we define the symmetrization of $T_{1},T_{2}$ as
$T_{1}\underline{\otimes}T_{2}:=\frac{1}{|\mathcal{P}_{p_{1},p_{2}}|}\sum_{\pi\in\mathcal{P}_{p_{1},p_{2}}}(T_{1})_{\pi^{-1}(1)}(T_{2})_{\pi^{-1}(2)}.$
In the particular case where $T_{1}=T_{2}=T$, we write $T^{\underline{\otimes}2}$.
Wigner tensors.
We consider $X=(X_{i})_{i\in{{\left[\left[N\right]\right]}}^{p}}\in\mathcal{E}^{N}_{p}=(\mathbb{C}^{N})^{\otimes p}$ be such that $X_{i}=X_{j}$ if $i\stackrel{{\scriptstyle p}}{{\sim}}j$ and the random variables $(X_{i})_{i{{\left[\left[N\right]\right]}}^{p}/\stackrel{{\scriptstyle p}}{{\sim}}}$ are independent, real,
$\mathbb{E}X_{i}=0\quad\quad\mathrm{and}\quad\quad\mathbb{E}X_{i}^{2}=\frac{p}{\mathcal{P}_{i}},$ (8)
where $\mathcal{P}_{i}$ is the number of elements in the equivalence class of $i$. We assume moreover that all the moments are bounded, that is
  1. (X1)
    For all integers $k\geq 2$, there exists a constant $c(k)>0$ such that for all integers $N\geq 1$ and $i\in{{\left[\left[N\right]\right]}}^{p}$:
    $\mathbb{E}|X_{i}|^{k}\leq c(k).$
Remark that we can equivalently write $\mathbb{E}X_{i}^{2}=\frac{1}{(p-1)!}\prod_{j=1}^{N}c_{j}(i)!,$ where $c_{j}(i)$ is the number of occurrences of $j$ in $i$. We also assume that for each integer $N$, the law of $X_{i}$ depends only on the equivalence class of $i\in{{\left[\left[N\right]\right]}}^{p}$ with respect to the action of $\mathrm{S}_{N}$. The law of $X_{i}$ may depend on $N$.
The main example is the Gaussian Orthogonal Tensor Ensemble (GOTE) where $X$ is Gaussian. We define the normalized Wigner tensor as
$\mathbf{W}_{N}:=\frac{X}{N^{\frac{p-1}{2}}}.$
(Note that for vectors, $p=1$, there is no scaling). The random tensor $\mathbf{W}^{N}$ is the tensor analog of the real Wigner matrices.
Remark 1.
For the convergence results which follow, we could assume that the entries have all variance $\frac{1}{(p-1)!}$ because as in the matrix case the terms of leading order in the proof only make appear the entries outside any diagonal, that is when $i_{1},\ldots,i_{p}$ are two by two distincts and then $\frac{p}{\mathcal{P}_{i}}=\frac{1}{(p-1)!}$.
Wishart tensors.
We define another main type of real symmetric random tensors. We begin with the even case which is more natural. Let $p\geq 2$ be an even integer. Let $\{x_{i,j_{1},\ldots,j_{p/2}}:i,j\geq 1\}$ an array of i.i.d. real random variables with
$\mathbb{E}x_{i,j_{1},\ldots,j_{p/2}}=0\quad\quad\mathrm{and}\quad\quad\mathbb{E}x_{i,j_{1},\ldots,j_{p/2}}^{2}=\frac{1}{\eta_{p}},$ (9)
where $\eta_{p}:=[(p/2)!]^{2/p}$. We assume moreover that all the moments are bounded, that is
  1. (X1’)
    For all integers $k\geq 2$, there exists a constant $c^{\prime}(k)>0$ such that for all integers $N\geq 1$ and $i\in{{\left[\left[N\right]\right]}}^{p}$:
    $\mathbb{E}|x_{i,j_{1},\ldots,j_{p/2}}|^{k}\leq c^{\prime}(k).$
Then write $x_{l}^{N}:=(x_{l,j_{1},\ldots,j_{p/2}},\ldots,x_{l,j_{1},\ldots,j_{p/2}})_{1\leq j_{1},\ldots,j_{p/2}\leq N}\in\mathcal{E}_{p/2}^{N}$. We will in the following forget the $N$ in this notation for lightness. For $p\geq 2$, the tensor
$x_{l}^{\underline{\otimes}2}\in(\mathbb{R}^{N})^{\otimes p}$
is a symmetric tensor introduced in the previous paragraph.
Then we define the Wishart tensor of size $k$ as
$\mathcal{W}_{N}^{k}:=\frac{x_{1}^{\underline{\otimes}2}+\ldots+x_{k}^{\underline{\otimes}2}}{N^{\frac{p}{2}}}.$
Now assume that $k=k_{N}$ goes to infinity with $N$. More precisely, we are going to assume that $\frac{k_{N}}{N^{p/2}}\rightarrow t\in(0,\infty)$. The random tensor $\mathcal{W}^{N}$ is a Wishart matrix in the particular case $p=2$. Another tensor analog of the real Wishart matrices could be a sum of rank one tensors $x_{1}^{\otimes p}+\ldots+x_{k}^{\otimes p}$, but its moments does not converge whatever the normalization.
If $p$ is odd we choose $p_{1},p_{2}\geq 1$ such that $p=p_{1}+p_{2}$, for instance $p_{1}=\frac{p+1}{2}=p_{2}+1$. Let $\{x_{i,j_{1},\ldots,j_{p_{1}}}:i,j\geq 1\}$ and $\{y_{i,j_{1},\ldots,j_{p_{2}}}:i,j\geq 1\}$ two arrays of i.i.d. real random variables with
$\mathbb{E}x_{i,j_{1},\ldots,j_{p_{1}}}=\mathbb{E}y_{i,j_{1},\ldots,j_{p_{2}}}=0\quad\quad\mathrm{and}\quad\quad\mathbb{E}x_{i,j_{1},\ldots,j_{p_{1}}}^{2}=\frac{1}{[p_{1}!]^{1/p}},\mathbb{E}y_{i,j_{1},\ldots,j_{p_{2}}}^{2}=\frac{1}{[p_{2}!]^{1/p}}.$ (10)
We denote $\eta_{p}:=[p_{1}!p_{2}!]^{1/p}$. We then construct the Wishart tensor exactly as before
$\mathcal{W}_{N}^{k}:=\frac{x_{1}\underline{\otimes}y_{1}+\ldots+x_{k}\underline{\otimes}y_{k}}{N^{\frac{p}{2}}}.$
Remark 2.
In both Wigner and Wishart cases, the convergence of the moments that we are going to prove in the following still holds if there are only $p$ finite moments, and not necessarily all of them. We may then weaken assumptions $(X1)$ and $(X1^{\prime})$. To prove that, it is then necessary to define $\widehat{x}=x\mathbf{1}_{x\leq C}$ and $\widetilde{x}=\widehat{x}-\mathbb{E}\widehat{x}$ and to control the difference between $\mathbf{W}$ and $\widetilde{\mathbf{W}}$, or $\mathcal{W}$ and $\widetilde{\mathcal{W}}$. We will not treat this case here and we will assume for simplicity that all the moments are bounded.
3.2 Semicircular in high order
The law playing the role of the Gaussian is this free framework is the semicircular of high order, whose even moments are given by the Fuss-Catalan numbers. In the case $p=2$ this law is the semi-circle law and we retrieve usual results of free probability. We define the Fuss-Catalan numbers as
$F_{p}(k)=\frac{1}{pk+1}\binom{pk+1}{k},$
which counts the number of $p$-melonic graphs (that is greater than a disjoint union of melons of order $p$ in our poset) with $k$ vertices, or also the number of non-crossing partitions of $\{1,\ldots,(p-1)k\}$ with blocks of size multiple of $p-1$. A bijection can be find back by the partition on the half edges looking on the tree obtained after merging the pair-associated vertices.
Definition 3 (High order semicircular law).
For $p\in\mathbb{N}$ we define $\mu_{p}$ the law having for moments
[…]
This measure is compactly supported on $[-\sqrt{p^{p}/(p-1)^{p-1}},\sqrt{p^{p}/(p-1)^{p-1}}]$.
Then the measure associated to a Wigner tensor (and in particular a Gaussian tensor of the GOTE) converges weakly in probability towards $\mu_{p}$ when the dimension goes to infinity. It relies on the two following Propositions [Theorem $1$ and Theorem $2$ in [11]], we will recall some ideas of the proof of the convergence of the moments.
Proposition 1.
For all $n\geq 0$ and all $p\geq 2$,
$\mathbb{E}[m_{n}(\mathcal{W}_{N})]=m_{n}(\mu_{p})+\mathcal{O}(1/N).$
Proposition 2.
For all $n\geq 0$ and all $p\geq 2$,
$\mathrm{Var}[m_{n}(\mathcal{W}_{N})]=\mathcal{O}(1/N^{2}).$
Ideas of the proof of Proposition 1.
In some words, the proof consists in classifying the $p$-regular combinatorial maps on $n$ vertices with given indices $j_{1},\ldots,j_{np/2}$ associated to each edge into three categories. We define an equivalence relation on the vertices where $x\sim y$ if $x$ and $y$ belong to $p$ edges with the $p$ same indices $j$ associated. We choose one canonical graph in each isomorphism class (two graphs are isomorphic if one is the other after acting by a permutation on ${{\left[\left[N\right]\right]}}$ for the indices on each edge), and we classify :
  • $\bullet$
    Category $1$ : each equivalence class contains exactly two vertices (in particular $n$ is even), and the number of distinct indices is maximal. We denote $\mathcal{M}_{n}$ the set of these graphs.
  • $\bullet$
    Category $2$ : at least one equivalence class contain one single vertex.
  • $\bullet$
    Category $3$ : all the other graphs.
As the graph is connected the maximal number of distinct indices for a graph in Category $1$ is $1+\frac{n(p-1)}{2}$ ($p$ for the first pair and $p-1$ for each one of the $\frac{n}{2}-1$ others, which is also the consequence of an Euler formula on the reduced tree as there is no cycle). Moreover, there are
$N(N-1)\ldots(N-\frac{n(p-1)}{2})[(p-1)!]^{n/2}=N^{1+\frac{n(p-1)}{2}}[(p-1)!]^{n/2}(1+\mathcal{O}(1/N)$
graphs in each isomorphism class. We claim that $\mathcal{M}_{n}$ is exactly the set of the melonic graphs with $n$ vertices with good indices, whose number is given by $F_{p}(n/2)$. As for a graph in $\mathcal{M}_{n}$, $\mathbb{E}X_{G}=(\mathbb{E}X_{1,\ldots,p}^{2})^{n/2}=\frac{1}{[(p-1)!]^{n/2}}$, then the contribution of these graphs is
$\frac{1}{N}\frac{1}{N^{\frac{n(p-1)}{2}}}\sum_{\mathcal{M}_{n}}N(N-1)\ldots(N-\frac{n(p-1)}{2})\times 1=F_{p}(n/2)+\mathcal{O}(1/N).$
Furthermore, by centering of the entries, the graphs in Category $2$ gives a zero contributon. The ones in Category $3$ gives a contribution $\mathcal{O}(1/N)$. That concludes the proof, see [11] for details. ∎
To finish with the semicircular we give a proof of Theorem 6. Let $(T_{i})_{i\geq 1}$ as in our assumptions of the free Central Limit Theorem.
Proof of Theorem 6.
Let $s_{k}:=\frac{1}{\sqrt{k}}\sum_{i=1}^{k}T_{i}$. We compute the limit of $\kappa_{n}(s_{k})$ for all $n$ as $k\rightarrow\infty$. By free independence, the mixed cumulants are zero, so we have
$\kappa_{n}(s_{k})=k^{-n/2}\sum_{i=1}^{n}\sum_{\mathfrak{b}\in\mathcal{B}_{n}}\kappa_{\mathfrak{b}}(T_{i}).$
Firstly, $\kappa_{1}(s_{k})=m_{1}(s_{k})=0$. Moreover, as $\kappa_{2}(T_{i})=m_{2}(T_{i})-\frac{p^{2}}{4}m_{1}(T_{i})^{2}=1$, then $\kappa_{2}(s_{k})=\frac{k}{k}=1$. Finally, if $n\geq 3$, then
$|\kappa_{n}(s_{k})|\leq\frac{1}{k^{\frac{n}{2}-1}}\mathrm{max}_{\mathfrak{b}\in\mathcal{B}_{n}}C(\mathfrak{b})$
which goes to zero. That gives the result. ∎
3.3 Free Poisson in high order
The other main law of the free world is the free Poisson law, also called the Marčenko-Pastur, which is the limiting measure of a Wishart matrix. We are going to show that the measure associated to a Wishart tensor converges in probability towards a higher order free Poisson law. Before describing this law and proving the convergence we may introduce some new notions.
[…]
#$\#NC_{q}(n)=F_{q+1}(n)=\frac{1}{(q+1)n+1}\binom{(q+1)n+1}{n}.$
Moreover, for $1\leq b\leq n$ we denote $NC_{q}^{b}(n)$ the non-crossing partitions of ${{\left[\left[qn\right]\right]}}$ having $b$ blocks, whose sizes are multiple of $q$. Then we denote
$F_{q}^{b}(n):=\#NC_{q}^{b}(n)$
the cardinal of such partitions.
Integer case.
For $q\in\mathbb{N}$, they have been counted, see [19], and their number is given by the Fuss-Narayana numbers, that is
$F_{q}^{b}(n)=\frac{1}{b}\binom{n-1}{b-1}\binom{qn}{b-1},$
which are a generalization of Narayana numbers.
Half-integer case.
Now if $q=q^{\prime}/2$ with $q\notin\mathbb{N},q^{\prime}\in\mathbb{N}$, then the blocks of a partition having size multiple of $q$ must have a size of length multiple of $q^{\prime}$ (as $q\mathbb{N}\cap\mathbb{N}=q^{\prime}\mathbb{N}$). Hence, as $qn\notin\mathbb{N}$ when $n$ is odd, we can write in the half-integer case
$F_{q}^{b}(n)=\left\{\begin{array}[]{ll}0&\mbox{if }n\mbox{ odd}\\ F_{2q}^{b}(n)=\frac{1}{b}\binom{n-1}{b-1}\binom{2qn}{b-1}&\mbox{if }n\mbox{ even.}\end{array}\right.$
Definition 4 (Free Poisson of order $p$).
For $p\geq 2$ even, $t\in(0,\infty)$ we define $\nu_{p,t}$ the law having for moments
[…]
This measure is compactly supported.
When $p$ is even, this law has for moments the Fuss-Narayana polynomials of parameters $\frac{p}{2},t$ and it it then a so called free Bessel law $\pi_{\frac{p}{2},t}$ as described in [8] where it is in particular proved that this measure has a compact support. When $p$ is odd, the odd moments are zero and the even moments are given by the Fuss-Narayana numbers of parameters $p,t$. This law is then the image by the map $z\mapsto z^{2}$ of a free Bessel law of parameters $p,t$.
The existence of this law is then contained in [8].
Case $t=1$.
A relevant point we mention here is the particular case $t=1$. We retrieve in this case the Fuss-Catalan numbers observed for the semicircular, with a different $p$. This means in particular that a semicircular law of order $p+1$ is the image of a free Poisson law of order $2p$ by the map $z\mapsto z^{2}$. In the matrix case, we retrieve that the semicircular (of order $2=1+1$) is the image by the map $z\mapsto z^{2}$ of a quarter-circular law (free Poisson - or Marčenko-Pastur as $t=t^{-1}=1$ of order $2=2\times 1$).
Case $t\rightarrow\infty$.
When the parameter $t$ goes to infinity, the reshaped high order free Poisson law $\frac{1}{\sqrt{t}}(\nu_{p,t}-\Delta_{t})$ tends to the high order semicircular. Here $\Delta_{t}$ is the probability measure with first cumulant equal to $t$ and others equal to 00. We postpone the proof of this claim to the following Section, see Remark 5.
We are now ready to focus on the main result of this section, that is the weak convergence in probability of the measure associated to $\mathcal{W}^{p,k}_{N}$ a $p$-order Wishart tensor such that $k_{N}/N^{p/2}\rightarrow t\in(0,\infty)$ towards $\nu_{p,t}$ as the dimension grows, given by Theorem 5. The proof of this theorem rely on the moments method. It is indeed sufficient to prove the two following propositions.
Proposition 3.
For all $n\geq 0$ and all $p\geq 2$, when $N\rightarrow\infty$,
$\mathbb{E}[m_{n}(\mathcal{W}_{N})]=\sum_{b=1}^{n}F^{b}_{p/2}(n)t^{b}+\mathcal{O}(1/N).$
Proposition 4.
For all $n\geq 0$ and all $p\geq 2$, when $N\rightarrow\infty$,
$\mathrm{Var}[m_{n}(\mathcal{W}_{N})]=\mathcal{O}(1/N^{2}).$
Remark 3.
We want to call attention on a tricky point. The most classical result about convergence of Wishart matrices towards Marčenko-Pastur law treat the case where $N/k_{N}\rightarrow\tau$ where as mentioned before the limiting Marčenko-Pastur law is equal to the free Poisson law of parameter $t=1/\tau$ dilated by a factor $\tau^{p/2}=\tau$. The usual Marčenko-Pastur law has also moments given by the Narayana polynomials due to the fact that
$\tau^{n}\sum_{b=1}^{n}\frac{1}{b}\binom{n-1}{b-1}\binom{n}{b-1}\frac{1}{\tau^{b}}=\sum_{r=0}^{n-1}\frac{1}{n-r}\binom{n-1}{r}\binom{n}{r+1}\tau^{r}=\sum_{r=0}^{n-1}\frac{1}{r+1}\binom{n-1}{r}\binom{n}{r}\tau^{r}.$
This is no more the case when $p>2$.
3.3.1 Proof of Propostion 3
The ideas of this proof follow the ones of Bai and Silverstein in [[5], Section 3] to prove classical convergence towards the Marčenko-Pastur law, adapted in our setting. Note first that the terms due to different choices of $\pi$ in the symmetrization do not interact as the $x_{i,j}$ are chosen i.i.d. As the computation will not depend on the pairing the normalization by $\mathcal{P}_{p_{1},p_{2}}$ will disappear with the sum on $\mathcal{P}_{p_{1},p_{2}}$. Hence we fix the canonical $\pi$ ($p_{1}$ first indices and $p_{2}$ last ones) in the rest of the proof and study the asymmetric tensor $(x_{1}\otimes y_{1}+\ldots+x_{k}\otimes y_{k})/N^{\frac{p}{2}}$.
The proof can start. We begin by defining a class of graphs and proving some lemmas about their combinatorics.
$\Delta$-graphs.
Let $i_{1},\ldots i_{n}$ be $n$ integers (not necessarily distinct) in ${{\left[\left[k\right]\right]}}$, and $j_{1},\ldots,j_{np}$ be $np$ integers (not necessarily distinct) in ${{\left[\left[N\right]\right]}}$. Denote $b$ the number of distinct elements in $\{i_{1},\ldots i_{n}\}$ and plot $i_{1},\ldots i_{n}$ on $b$ parallel lines $I_{1},\ldots,I_{b}$ ordered by the first element seen in $i_{1},\ldots i_{n}$. Each line $I_{u}$ has one vertex and we say that it is a vertex of type $I$. Then, plot $b$ parallel lines $J_{1},\ldots,J_{b}$ such that $J_{u}$ is between $I_{u}$ and $I_{u+1}$. Now, recursively for $1\leq v\leq n$, if $i_{v}$ is on the line $I_{u}$, then plot the not already plotted $j_{(v-1)p+1},\ldots,j_{vp}$ on $I_{u}$ and draw $p$ edges from $i_{v}$ to $j_{(v-1)p+1},\ldots,j_{vp}$ (that can be multiple edges if these $j$ are not distinct). Denote $1+r$ the number of distinct elements in $\{j_{1},\ldots,j_{np}\}$, that is the number of vertices of type $J$. Such a graph always has $kp$ edges. A $\Delta$-graph is a graph constructed like that, with the two additional conditions :
  1. 1.
    each vertex of type $J$ has valence multiple of $2$ (belongs to $2m$ edges),
  2. 2.
    the graph is connected.
An example is given in Figure 4. If $p$ is even, for $1\leq v\leq n$, we call the $(p/2)$-tuples of edges between $i_{v}$ and $(j_{(v-1)p+1},\ldots,j_{(v-1)p+\frac{p}{2}})$, respectively $(j_{(v-1)p+\frac{p}{2}+1},\ldots,j_{vp})$, the two $v$-multiedges denoted $e_{v}(1)$ and $e_{v}(2)$. We say that two multiedges $e_{u}(i)$ and $e_{v}(i^{\prime})$ are paired if one can pair them into $p/2$ coincident edges, that is $i_{u}=i_{v}$ and there exists $\sigma\in S_{p/2}$ such that for all $1\leq r\leq\frac{p}{2},j_{(u-1)p+(i-1)\frac{p}{2}+r}=j_{(v-1)p+(i^{\prime}-1)\frac{p}{2}+\sigma(r)}$. When $p$ is odd, we do the same construction with $e_{v}(1)$ the $p_{1}$ first edges and $e_{v}(2)$ the $p_{2}$ last ones. In the even case, it can be $u=v$ and $i=1,i^{\prime}=2$, but not in the odd case.
[…]
Figure 4: A $\Delta$-graph ($p=4,n=3,b=2,r+1=5$).
Two graphs are said isomorphic if there are the same up to a permutation on $(1,\ldots,k)$ and a permutation on $(1,\ldots,N)$. For each isomorphism class, there is only one canonical graph satisfying $i_{1}=j_{1}=1$, $i_{u+1}\leq\mathrm{max}\{i_{1},\ldots,i_{u}\}+1$ and $j_{u+1}\leq\mathrm{max}\{i_{1},\ldots,i_{u}\}+1$. The set of canonical $\Delta$-graphs with $b$ vertices of type $I$ and $1+r$ vertices of type $J$ is denoted $\Delta(n,b,r)$. We then classify the $\Delta(n,b,r)$-graphs into $3$ categories :
  • $\bullet$
    Category 1 : $\Delta(n,b,r)$-graphs in which there is no cycle, each edge coincides with one and only one other edge, and the multiedges are paired (two by two). If we glue the double edges, the resulting graph is a tree with $kp/2$ edges. Hence, $\frac{kp}{2}=b+r$ and we can juste denote these graphs the set of $\Delta_{1}(n,b)$-graphs as $r$ can be forgotten.
  • $\bullet$
    Category 2 : $\Delta(n,b,r)$-graphs that contain at least one non-paired multiedge. We denote them the $\Delta_{2}(n,b,r)$.
  • $\bullet$
    Category 3 : $\Delta(n,b,r)$-graphs which do not belong to $\Delta_{1}(n,b)$ or $\Delta_{2}(n,b,r)$. We denote them the $\Delta_{3}(n,b,r)$-graphs.
An example of $\Delta_{1}(n,b)$-graph is given in Figure 5.
[…]
Figure 5: A $\Delta_{1}(n,b)$-graph.
We can now prove some results about their combinatorics.
Lemma 3.
For given $n,b,r$, the number of graphs in the isomorphism class of $G\in\Delta(n,b,r)$ is
$k(k-1)\ldots(k-b+1)N(N-1)\ldots(N-r)c_{p}^{G}=k^{b}N^{r+1}c^{G}_{p}(1+\mathcal{O}(N^{-1}).$
where $c^{G}_{p}$ is a constant depending only on the shape of $G$, with in particular if $G\in\Delta_{2}(n,b)$,
$c^{G}_{p}=\eta_{p}^{np/2}.$
Proof.
The factor $\eta_{p}^{np/2}$ comes from the choice of the pairing of the $j$ in each pair of multiedges. One must distinguish the even and the odd cases. In the even case, we have $c_{p}^{G}=[(p/2)!]^{n}$. In the odd case, $n$ must be even as we cannot pair multiedges from a same vertex of type I, so $c_{p}^{G}=[p_{1}!p_{2}!]^{n/2}$. In both cases, that is equal to $\eta_{p}^{np/2}$. The rest is clear. ∎
Lemma 4.
The total number of noncoincident vertices of a $\Delta_{3}(n,b,r)$-graph is at most $np/2$.
Proof.
Let $G$ be a graph in $\Delta_{3}(n,b,r)$. Since $G$ is not in $\Delta_{2}$, it has no single edge, so the number of noncoincident edges is at most $np/2$. As any $\Delta$-graph is connected, the number of noncoincident vertices is at most the number of noncoincident edges plus one with equality if and only if the noncoincident graph is a tree. But $G$ is not in $\Delta_{1}$, so if it has exactly $np/2$ noncoincident vertices, then it must contain a cycle. Hence, in all cases the number of noncoincident vertices is at most $np/2$. ∎
The last very important result is to count the number of $\Delta_{1}(n,b)$-graphs.
Lemma 5.
For $n,b$, the number of $\Delta_{1}(n,b)$-graphs is
$F^{b}_{p/2}(n).$
Proof.
This lemma stands on a bijection between $\Delta_{1}(n,b)$-graphs and non-crossing partitions of ${{\left[\left[np/2\right]\right]}}$ with $b$ block,s whose sizes are multiple of $\frac{p}{2}$. For a $\Delta_{1}(n,r)$-graph $G$ we glue the coincident edges, it remains $\frac{np}{2}$ edges which we label by their order of visit in the depth-first search starting from $x_{1}$. Then we associate to this graph the (non-crossing) partition on ${{\left[\left[np/2\right]\right]}}$ with $b$ blocks where $j_{1},j_{2}$ belong to the same block if and only if the vertex of type $I$ belonging to the edges labeled $j_{1}$ and $j_{2}$ is the same. This bijection gives also easily the way to reconstruct a $\Delta_{1}(n,r)$-graph from a partition in $\mathrm{NC}^{b}_{p/2}(n)$. ∎
We are now ready to prove the convergence of the moments.
Proof of Proposition 3..
A first calculus gives :
$\mathbb{E}[m_{n}(\mathcal{W}_{N})]$ $=\frac{1}{N}\frac{1}{k^{n}}\sum_{\mathfrak{b}\in\mathcal{B}_{n}}\mathbb{E}\mathfrak{b}(x_{1}^{\underline{\otimes}2}+\ldots+x_{k}^{\underline{\otimes}2})$
$=\frac{1}{N}\frac{1}{N^{np/2}}\sum_{\mathfrak{b}=(V,E)\in\mathcal{B}_{n}}\sum_{1\leq i_{1},\ldots,i_{n}\leq k}\sum_{1\leq j_{1},\ldots,j_{np/2}\leq N}\prod_{a=1}^{n}\mathbb{E}(x_{i_{a}})_{j_{(\delta v_{a})_{1}}}\ldots(x_{i_{a}})_{j_{(\delta v_{a})_{p}}}.$
Two terms are equal if the corresponding graphs are isomorphic. Therefore by Lemma 3,
$\mathbb{E}[m_{n}(\mathcal{W}_{N})]=\frac{1}{N}\frac{1}{k^{n}}\sum_{G\in\Delta(n,b,r)}k(k-1)\ldots(k-b+1)N(N-1)\ldots(N-r)c_{p}^{G}\mathbb{E}\mathbf{x}_{G},$
where the summation is over all canonical $\Delta(n,b,r)$-graphs. The summation is into three parts according to the type of the graph $\Delta_{a}(n,b,r)$ for $a=1,2$ or $3$. Firstly, if $G$ is a $\Delta_{2}(n,b,r)$-graph then $\mathbb{E}\mathbf{x}_{G}=0$ as the $x_{i}$ are centered and there exists at least one non-paired multiedge. That is,
$S_{2}=\frac{1}{N^{1+np/2}}\sum_{G\in\Delta_{2}(n,b,r)}k(k-1)\ldots(k-b+1)N(N-1)\ldots(N-r)c_{p}^{G}\mathbb{E}\mathbf{x}_{G}=0.$
Secondly, if $G$ is a $\Delta_{3}(n,b,r)$-graph then $b+r<np/2$ by Lemma 4 and as $\mathbb{E}\mathbf{x}_{G}\leq c^{\prime}(n)$ then
$S_{3}=\frac{1}{N^{1+np/2}}\sum_{G\in\Delta_{3}(n,b,r)}k(k-1)\ldots(k-b+1)N(N-1)\ldots(N-r)c^{G}_{p}\mathbb{E}\mathbf{x}_{G}=\mathcal{O}(N^{-1}).$
Finally, if $G$ is a $\Delta_{3}(n,b)$-graph then $\mathbb{E}\mathbf{x}_{G}=(\mathbb{E}[(x_{1})_{1}^{2}])^{np/2}=1/\eta_{p}^{np/2}$. So by Lemma 5,
$S_{1}$ $=\frac{1}{N^{1+np/2}}\sum_{G\in\Delta_{1}(n,b)}k(k-1)\ldots(k-b+1)N(N-1)\ldots(N-r)\eta_{p}^{np/2}\mathbb{E}\mathbf{x}_{G}$
$=\frac{1}{N^{1+np/2}}\sum_{b=1}^{n}F_{p/2}^{b}(n)k^{b}N^{1+\frac{np}{2}-b}+\mathcal{O}(N^{-1})$
$=\sum_{b=0}^{n-1}F_{p/2}^{b}(n)t_{N}^{b}+\mathcal{O}(N^{-1}),$
where $t_{N}=k_{N}/N^{p/2}\rightarrow t$. This concludes the proof. ∎
3.3.2 Proof of Proposition 4
We only need to show that $\mathrm{Var}[m_{n}(\mathcal{W}_{N})]$ is summable for all fixed $n$.
Proof of Proposition 4.
We now have
$\mathrm{Var}[m_{n}(\mathcal{W}_{N})]=\frac{1}{N^{2}}\frac{1}{N^{np}}\sum_{G_{1},G_{2}\in\Delta}[\mathbb{E}\mathbf{x}_{G_{1}}\mathbf{x}_{G_{2}}-\mathbb{E}\mathbf{x}_{G_{2}}\mathbb{E}\mathbf{x}_{G_{1}}].$
Firstly, if $G_{1}$ has no coincident edge with $G_{2}$, then by independence
$\mathbb{E}\mathbf{x}_{G_{1}}\mathbf{x}_{G_{2}}-\mathbb{E}\mathbf{x}_{G_{2}}\mathbb{E}\mathbf{x}_{G_{1}}=0.$
Moreover, if $G_{1}\cup G_{2}$ has an overall single edge, then we have also
$\mathbb{E}\mathbf{x}_{G_{1}}\mathbf{x}_{G_{2}}=\mathbb{E}\mathbf{x}_{G_{2}}\mathbb{E}\mathbf{x}_{G_{1}}=0.$
Hence, we can assume that $G_{1}$ has a coincident edge with $G_{2}$ and contains $G=G_{1}\cup G_{2}$ has no single edge. If the graph of noncoincident edges has a cycle, then the noncoincident edges of $G$ are not more than $np$. In the other case, if the graph of noncoincident edges has no cycle, then the number of noncoincident vertices is also not larger than $np$ due to the following reason. At least one edge must have coincidence greater than $3$, because otherwise we find an edge of coincidence $1$ in both $G_{1}$ and $G_{2}$, and then a second one as the number of edges is even, so that gives a cycle in $G$. Thus, either this edge has coincidence greater than $4$, either it is equal to $3$ and there is another one of coincidence greater than $3$ since again the number of edges is even. That gives that the number of noncoincident vertices is not larger than $(np-1)+1=np$. Consequently, we get
$\mathrm{Var}[m_{n}(\mathcal{W}_{N})]\leq\frac{1}{N^{2}}2c^{\prime}(n)K_{n},$
$K_{n}$ depending only on $n$. The proof is complete. ∎
4 Tensorial free convolution
The frame of this part is largely inspired by the lecture notes of Roland Speicher [31]. For this section we fix $p\geq 2$ even integer.
4.1 Free convolution of compactly supported measures
Let $\mu$ be a probability measure on $\mathbb{R}$ with compact support, that is there exists $M>0$ such that $\mu[-M,M]=1$. Then the moments $(m_{n}(\mu))_{n\in\mathbb{N}}$ of $\mu$,
  • (i)
    are all finite,
  • (ii)
    are exponentially bounded with constant $M$, i.e. for all $n\in\mathbb{N}$,
    $|m_{n}(\mu)|\leq M^{n},$
  • (iii)
    determine uniquely the probability measure $\mu$.
Proposition 5.
Let $a$ be a distribution on the trace maps with moments $(m_{n}(a))_{n\in\mathbb{N}}$ and free cumulants $(\kappa_{n}(a))_{n\in\mathbb{N}}$. Then the following statements are equivalent :
  • (i)
    the sequence $(m_{n}(a))_{n\in\mathbb{N}}$ is exponentially bounded,
  • (ii)
    the sequence $(\kappa_{n}(a))_{n\in\mathbb{N}}$ is exponentially bounded.
Proof.
$\boxed{(ii)\Rightarrow(i)}$ Assume that $\kappa_{n}(a)\leq M^{n}$ for all $n$. Then for $n\geq 0$,
$|m_{n}(a)|\leq\sum_{b=1}^{n}\sum_{\pi_{n}=(\pi_{i_{1}},\ldots,\pi_{i_{b}})\in\mathrm{NC}_{p}^{b}(n)}\underbrace{\sum_{\mathfrak{b}_{1},\ldots,\mathfrak{b}_{b}\in\mathcal{B}_{i_{1}}\times\ldots\times\mathcal{B}_{i_{b}}}|\kappa_{\mathfrak{b}_{1}}(a)\ldots\kappa_{\mathfrak{b}_{1}}(a)|}_{\leq M^{n}}\leq F_{p}(n)M^{n}\leq(2^{p}M)^{n}.$
$\boxed{(i)\Rightarrow(ii)}$ For $n\geq 0$, we have that
$\kappa_{\mathfrak{b}}(a)=\sum_{\mathfrak{b}^{\prime}\leq\mathfrak{b}}\mathrm{Moeb}(\mathfrak{b}^{\prime},\mathfrak{b})\mathfrak{b}^{\prime}(a),$
where $\mathrm{Moeb}(\mathfrak{b}^{\prime},\mathfrak{b})$ is the Moebius function associated to our poset. It is defined as follows : for any $\mathfrak{b}$, $\mathrm{Moeb}(\mathfrak{b},\mathfrak{b})=1$ and for $\mathfrak{b}^{\prime}<\mathfrak{b}$, $\mathrm{Moeb}(\mathfrak{b}^{\prime},\mathfrak{b})=-\sum_{\mathfrak{b}^{\prime}\leq\mathfrak{b}^{\prime\prime}<\mathfrak{b}}\mathrm{Moeb}(\mathfrak{b}^{\prime},\mathfrak{b}^{\prime\prime})$. Note that if $\mathfrak{b}\in\mathcal{B}_{n}$ and $\mathfrak{b}^{\prime}\leq\mathfrak{b}$, then $|\mathrm{Moeb}(\mathfrak{b}^{\prime},\mathfrak{b})|\leq F_{p}(n)$. Now assume that $m_{n}(a)\leq M^{n}$ for all $n$. Then for $n\geq 0$,
$|\kappa_{n}(a)|\leq F_{p}(n)\sum_{b=1}^{n}\sum_{\pi_{n}=(\pi_{i_{1}},\ldots,\pi_{i_{b}})\in\mathrm{NC}_{p}^{b}(n)}\underbrace{\sum_{\mathfrak{b}_{1},\ldots,\mathfrak{b}_{b}\in\mathcal{B}_{i_{1}}\times\ldots\times\mathcal{B}_{i_{b}}}|\kappa_{\mathfrak{b}_{1}}(a)\ldots\kappa_{\mathfrak{b}_{1}}(a)|}_{\leq M^{n}}\leq(4^{p}M)^{n}.$
Proposition 6.
Let $\mu$ and $\nu$ be two compactly supported probability measures such that
  • $\bullet$
    there exists two distributions $a$ and $b$ such that $\mu_{a}=\mu$ and $\mu_{b}=\nu$,
  • $\bullet$
    $a$ and $b$ are free.
The moments of $a+b$ are exponentially bounded and hence determine uniquely a compactly supported probability measure, denoted $\mu\oplus_{p}\nu$.
Proof.
By free independence of $a$ and $b$, we know that for all $n\geq 1$,
$\kappa_{n}(a+b)=\kappa_{n}(a)+\kappa_{n}(b).$
Hence, thanks to Proposition 5, we have that the moments of $a+b$ are exponentially bounded. This gives the result. ∎
Remark 4.
It does not depend on the choice of $a$ and $b$ because they are free so the free cumulants, summing over all maps with $n$ vertices, are whatever additive. However, not all distributions will correspond to a real tensor (eventually limiting) distribution.
Definition 5.
The probability measure $\mu\oplus_{p}\nu$ is called the tensorial free convolution of $\mu$ and $\nu$.
When $p=2$ we retrieve the usual free convolution. Now, let $\mu,\nu,\pi$ be compactly supported probability measures. Then the operation $\oplus_{p}$ has the following properties :
  1. (N)
    Neutral element : $\delta_{0}\oplus_{p}\mu=\mu$,
  2. (C)
    Commutativity : $\mu\oplus_{p}\nu=\nu\oplus_{p}\mu$,
  3. (A)
    Associativity : $\mu\oplus_{p}(\nu\oplus_{p}\pi)=(\mu\oplus_{p}\nu)\oplus_{p}\pi$.
Here $\delta_{0}$ is the measure having all zero moments and free cumulants for $n\geq 1$, associated to the distribution on $p$-regular maps equal to zero on all the non-empty maps associated. More generally, for $p$ even, the distribution $t.1_{p}$ is the distribution having for free cumulants
$\kappa_{\mathfrak{b}}(t.1_{p})=\frac{t}{(p-1)(p-3)\ldots 1}\mbox{ if }\mathfrak{b}\mbox{ is a bouquet map and }0\mbox{ otherwise.}$.
It is the identity appearing in [12], but symmetrized, and multiplied by a factor $t$. It has an associated probability measure on $\mathbb{R}$, denoted $\Delta_{t}$ having for free cumulants
$\kappa_{n}(\Delta_{t})=t\mbox{ if }n=1\mbox{ and }0\mbox{ if }n\geq 2,$
and for moments
$m_{n}(\Delta_{t})=F_{p/2}(n)t^{n}.$
This follows from Theorem 1, knowing that the functional relation satisfied by the $p$-Fuss-Catalan generative function is $f(z)=1+zf(z)^{p}$. This means in particular that the identity $1_{4}$ of order $4$ has for moments the Catalan numbers. Importantly, these distributions $t.1_{p}$ are free from anything as we proved in [[12], Lemma 7].
4.2 Basic examples for the semicircular and free Poisson
Firstly, we note that the analytic moment-cumulant formula given by Theorem 1 allows to retrieve the sequence of free cumulants from the sequence of moments and reciprocally with less effort than the combinatorial moment-cumulant formula on the poset. That is what gives the two following lemmas.
Lemma 6.
For $n\geq 1$, the free cumulants of the semicircular law of order $p$ are given by
$\kappa_{n}=\mathbf{1}_{n=2}.$
Proof.
Assume that $\pi$ is such that $C_{\pi}(z)=1+z^{2}$. Then,
$M_{\pi}(z)=1+z^{2}M_{\pi}(z)^{p}.$
Hence a simple recursive argument gives that the odd moments are zero, so $M_{\pi}(z)=f(z^{2})$ and we retrieve the functional equation of the Fuss-Catalan numbers $f(z)=1+zf(z)^{p}$. That gives that $\pi=\mu_{p}$. ∎
Lemma 7.
For $n\geq 1$, the free cumulants of a free Poisson law of order $p$ and parameter $t$ are given by
$\kappa_{n}=t$
if $p$ is even, and
$\kappa_{n}=\mathbf{1}_{n\text{ even }}t$
if $p$ is odd.
Proof.
We treat the even case, the odd case follows. Assume that $\pi$ has all non trivial free cumulants equal to $t$. Then for $|z|<1$,
$C_{\pi}(z)=1+\frac{zt}{1-z}.$
That gives the following relation for $M_{\pi}(z)$
$M_{\pi}(z)=1+zM_{\pi}^{p/2}(z)(M_{\pi}(z)+t-1),$
which is the equation satisfied by the generative function of the Fuss-Narayana numbers, see [8]. That means that $\pi=\nu_{p,t}$. ∎
Remark 5.
When the parameter $t$ goes to infinity, the reshaped free Poisson law $\frac{1}{\sqrt{t}}(\nu_{p,t}-\Delta_{t})$ tends to a semicircular. Indeed, since $1_{p}$ is free from anything as we mentioned just before,
$\kappa_{n}(b_{p,t}-t.1_{p})=0\mbox{ if }n=1\mbox{ and }t\mbox{ if }n\geq 2.$
Then, we have that
$\kappa_{2}(\frac{1}{\sqrt{t}}(b_{p,t}-t.1_{p}))=1,$
and for all $n\geq 3$, when $t\rightarrow\infty$,
$\kappa_{n}(\frac{1}{\sqrt{t}}(b_{p,t}-t.1_{p}))=t^{1-\frac{n}{2}}\rightarrow 0.$
That gives the result.
We give in the following the basic examples of tensorial free convolution for high order semicircular and free Poisson laws.
Lemma 8.
The free convolution of two freely independent semicirculars of order $p$ is a semicircular of order $p$ dilated by a factor $\sqrt{2}$. That is,
$\mu_{p}\oplus_{p}\mu_{p}=\mu_{p}^{(\sqrt{2})}.$
Proof.
One can just write
$m(z):=M_{\mu_{p}\oplus_{p}\mu_{p}}(z)=C_{\mu_{p}}(sm(z)^{p/2})+C_{\mu_{p}}(sm(z)^{p/2})-1=1+2z^{2}m(z)^{p}.$
That gives the result. ∎
Lemma 9.
The free convolution of two freely independent free Poisson of order $p$ and parameters $t$ and $t^{\prime}$ is a free Poisson of order $p$ and parameter $t+t^{\prime}$. That is,
$\nu_{p,t}\oplus_{p}\nu_{p,t^{\prime}}=\nu_{p,t+t^{\prime}}.$
Proof.
One can just write for $|z|<1$,
$m(z):=M_{\nu_{p,t}\oplus_{p}\nu_{p,t^{\prime}}}(z)=C_{\nu_{p,t}}(sm(z)^{p/2})+C_{\nu_{p,t^{\prime}}}(sm(z)^{p/2})-1=1+\sum_{n\geq 1}(t+t^{\prime})(zm(z)^{p/2})^{n}.$
Again, that gives the result. ∎
4.3 $R$-transform
Definition 6 ($R$-transform).
For $\mu$ a probability measure, its $R$-transform is the formal power series
$R_{\mu}(z)=\frac{C_{\mu}(z)-1}{z}=\sum_{n=1}^{\infty}\kappa_{n}(\mu)z^{n-1}.$
If $\mu$ is compactly supported, $R_{\mu}(z)$ converges for $|z|$ sufficiently small and we then have for any $\mu,\nu$ compactly supported probability measures
$R_{\mu\oplus_{p}\nu}(z)=R_{\mu}(z)+R_{\nu}(z),$
for $|z|$ sufficiently small.
We now define other version of transforms that could be more useful in the tensorial setting. We first recall some already known properties about the Cauchy transform of a measure, which we will not prove gain, see [31, 15].
Proposition 7.
[…]
  1. 1.
    $g_{\mu}:\mathbb{C}^{+}\mapsto\mathbb{C}^{-}$ is analytic on $\mathbb{C}^{+}$, and satisfies
    $\underset{y\rightarrow\infty}{\mathrm{lim}}iyg_{\mu}(iy)=1,$
  2. 2.
    any probability measure can be recovered from its Cauchy tranform via the inversion formula
    $\underset{\epsilon\rightarrow 0}{\mathrm{lim}}\frac{-1}{\pi}\int_{x}^{y}\mathrm{Im}[g_{\mu}(t+i\epsilon)]=\mu((x,y))+\frac{\mu(\{a,b\}}{2},$
  3. 3.
    if $\mu$ is compactly supported on $[-M,M]$ for some $R>0$, then $g_{\mu}$ has a power series expansion as follows
    $g_{\mu}(z):=\sum_{n=0}^{\infty}\frac{m_{n}(\mu)}{z^{n+1}}\mbox{ for all }z\in\mathbb{C}^{+}\mbox{ with }|z|>r.$
Remark 6.
For general tensors there is an associated notion of resolvent $g(T)$ for tensors, proposed by Gurau, and defined as an integral over $\mathbb{R}^{N}$, see [21, 11]. It has the formal expansion given by $g(T)=\sum_{n=0}^{\infty}\frac{m_{n}(\mu)}{z^{n+1}}$.
For ease of notataions we forgot the measure $\mu$ as index of all the moments, free cumulants and all the concernd formal series. The last point of the previous Proposition shows that $g(z)$ is a version of the moment series $M(z)=\sum_{n\geq 0}m_{n}z^{n}$, namely
$g(z)=\frac{1}{z}M\left(\frac{1}{z}\right).$
Theorem 1 implies that
$M\left(\frac{1}{z}\right)=C\left(\frac{1}{z}M\left(\frac{1}{z}\right)^{p/2}\right),$
hence denoting $G(z):=\frac{1}{z}M\left(\frac{1}{z}\right)^{p/2}$, we have
$C(G(z))^{p/2}=zG(z).$
Now we define the formal Laurent series $K(z):=\frac{C_{T}(z)^{p/2}}{z}$ and the previous equation gives
$K(G(z))=z,$ (11)
hence also
$G(K(z))=z.$ (12)
Since $K(z)$ has a pole $\frac{1}{z}$ we split this off and write
$K(z)=\frac{1}{z}+Q(z)\quad\mbox{where}\quad Q(z):=\sum_{n=1}^{\infty}\sum_{i_{1},\ldots,i_{p/2}\atop i_{1}+\ldots+i_{p/2}=n}\kappa_{i_{1}}\ldots\kappa_{i_{p/2}}z^{n-1}.$
Definition 7 ($Q$-transform).
For $\mu$ a probability measure and $p$ even integer, we define its tensorial $Q$-transform as the formal power series
$Q_{\mu}(z)=\frac{C_{\mu}(z)^{p/2}-1}{z}.$
When $\mu$ is no more compactly supported, one should study analytic properties of $G$ to deduce the ones of $Q$. We know that $G(z)=z^{\frac{p}{2}-1}g(z)^{p/2}$ and the analytic properties are well known (injective and has an inverse sufficiently far from 00). We may then define the subordination functions and try to study the additive free convolution of two general tensors. This will be part of further works.

### v2

@@
 1 Introduction
-This work follows [12] and we refer to that previous study for a more detailed introduction to freeness for tensors.
-In this paper, we further develop and expand upon this notion of tensorial freeness.
 The foundational and profoundly deep notion of freeness was introduced several decades ago by Dan-Virgil Voiculescu [33, 35].
@@
 This field is currently a vibrant area of research, driven by numerous applications in physics and computer science [2, 23, 10, 4, 22].
-Hence, in this paper, we aim to establish analogous results in the framework of tensorial freeness.
+In this paper, we aim to establish analogous results within the framework of tensorial freeness.
+Specifically, we proposed a notion of freeness for tensors in [12], to which we refer for more detailed introduction.
+Here, we continue developing the foundations of tensorial free probability theory, with a particular focus on the concept of tensorial free convolution.
 The principal contributions of this work are the following :
@@
 The distribution of a collection of tensors $\mathcal{A}=\{A_{1},\ldots,A_{n}\}$ of possibly different order, $A_{k}\in\mathcal{E}^{N}_{p_{k}}$ is the collection of all trace maps $\mathfrak{b}((T_{v})_{v\in V})$ with $T_{v}\in\mathcal{A}$ and $\mathfrak{b}$ with compatible degrees.
-In particular, if $T\in\mathcal{E}^{N}_{p}$, the distribution of $T$ is the collection of trace invariants for $p$-regular trace maps, that is all vertices of the map belong to exactly $p$-edges.
+In particular, if $T\in\mathcal{E}^{N}_{p}$, the distribution of $T$ is the collection of trace invariants for $p$-regular trace maps, that is all vertices of the map belong to exactly $p$ edges.
 We introduce now informally three types of $p$-regular trace maps useful for the rest of the introduction.
@@
 If $p$ is even, a bouquet is a map with one vertex and $p/2$ self-loops, and a multicycle is a map with $n\geq 1$ vertices $p/2$ edges between the $i$-th and the $i+1$-th ones for all $1\leq i\leq n$ ($n+1=1$).
-For $p$ odd, an odd multicycle is a map with $2n$ vertices and $(p+1)/2$ edges between the $12i-$ 1-th and the $2i$-th ones, $(p-1)/2$ edges between the $2i$-th and the $12i+$ 1-th ones, for all $1\leq i\leq n$ ($12n+1=$ 1).
+For $p$ odd, a variant of an odd multicycle is presented in Section 2.1.
 Non-crossing poset.
@@
 By transitivity, this gives a poset on the set of trace maps with a given number of vertices and a given sequence of degrees.
-The formal definitions are recall in Section 2 and an example is given in Figure 2.
+The formal definitions are recalled in Section 2 and an example is given in Figure 2.
 By Moebius inversion, it is possible to define free cumulants associated to this poset characterized by the implicit relation
@@
 1.1 Main results
-For a tensor $T$ of order $p$, and $n\geq 1$, we define $m_{n}(T)$, the $n$-th moment of $T$, as the sum over all the $p$-regular connected rooted trace maps with $n$ vertices of the trace invariants where all vertices are decorated by $T$.
-Similarly, $\kappa_{n}(T)$ the $n$-th free cumulant of $T$ is the sum over all the $p$-regular connected trace maps with $n$ vertices of free cumulant of the trace map where all vertices are decorated by $T$.
+Let $\mathcal{B}^{(p)}_{n}$ be the set of connected rooted $p$-regular trace maps with $n$ (unlabeled) vertices.
+For a tensor $T$ of order $p$, and $n\geq 1$, we define $m_{n}(T)$, the $n$-th moment of $T$, as the sum of $\mathfrak{b}(T,\ldots,T)$ over all $\mathfrak{b}\in\mathcal{B}^{(p)}_{n}$, where all vertices are decorated by $T$.
+Similarly, $\kappa_{n}(T)$ the $n$-th free cumulant of $T$ is the sum of $\kappa_{\mathfrak{b}}(T,\ldots,T)$ over all $\mathfrak{b}\in\mathcal{B}^{(p)}_{n}$.
 We set $m_{0}(T)=\kappa_{0}(T)=1$.
@@
 Theorem 1 (Analytic moment-cumulant formula).
-We have as functionnal relation in $\mathbb{C}{{\left[\left[z\right]\right]}}$,
+We have as functional relation in $\mathbb{C}{{\left[\left[z\right]\right]}}$,
 $M_{T}(z)=C_{T}\left(zM_{T}(z)^{p/2}\right).$
@@
 […]
-We have not yet succeeded to give a direct argument of the existence of this measure by showing the positivity of these moments, nevertheless the following results of convergence can always be given in terms of convergence of the moment and vanishing of the variance instead of in terms of weak convergence of the measure.
-If $a$ is a distribution on the $p$-regular combinatorial maps (that is $a:\mathcal{M}_{0}\mapsto\mathbb{R}$ with $a(\emptyset)=1$), we can speak of its moments $m_{n}(a)$ as the sum of $a(\mathfrak{b})$ over all the $p$-regular connected maps with $n$ vertices, and we denote $\mu_{a}$ as the measure on $\mathbb{R}$ having $(m_{n}(a))_{n\geq 0}$ as moments, if it exists.
+A direct proof of the existence of this measure is still missing.
+If $a$ is a distribution on the $p$-regular trace maps (that is $a:\cup_{n}\mathcal{B}^{(p)}_{n}\mapsto\mathbb{R}$ satisfying multilinearity, morphism property and $a(\emptyset)=1$), we may speak of its moments $m_{n}(a)$ as the sum of $a(\mathfrak{b})$ over all $\mathfrak{b}\in\mathcal{B}^{(p)}_{n}$, and we denote $\mu_{a}$ as the measure on $\mathbb{R}$ having $(m_{n}(a))_{n\geq 0}$ as moments, if it exists.
+By linearity we can equivalently define $a$ by giving $a(\kappa_{\mathfrak{b}})$.
 High order laws.
@@
 $a_{p}(\kappa_{\mathfrak{b}})=\frac{1}{(p-1)!}\mbox{ if }\mathfrak{b}\mbox{ is a melon and }0\mbox{ otherwise,}$,
-    if $p$ is even,
-    $b_{p,t}(\kappa_{\mathfrak{b}})=\frac{t}{[(p/2)!]^{n}}\mbox{ if }\mathfrak{b}\mbox{ is a multicycle of length }n\mbox{ and }0\mbox{ otherwise,}$,
-    if $p$ is odd,
-    $b_{p,t}(\kappa_{\mathfrak{b}})=\frac{t}{[(p+1/2)!(p-1/2)!]^{n}}\mbox{ if }\mathfrak{b}\mbox{ is an odd multicycle of length }2n\mbox{ and }0\mbox{ otherwise.}$.
+$b_{2p,t}(\kappa_{\mathfrak{b}})=\frac{t}{(p!)^{n}}\mbox{ if }\mathfrak{b}\mbox{ is a multicycle of length }n\mbox{ and }0\mbox{ otherwise,}$,
+Note that this is equivalent to $a_{p}(\mathfrak{b})=\mathbf{1}_{\mathfrak{b}\text{ melonic}}$.
+Moreover, the case of $p$ odd for $b$ is treated in Section 3.1 but not stated here for ease of notation.
 Theorem 2 (Semicircular law in high order).
@@
 and for free cumulants
-$\kappa_{2}(\mu_{p})=1\mbox{ if $n=0$ or $2$}\quad\quad\mathrm{and}\quad\quad\kappa_{n}(\mu_{p})=0\mbox{ otherwise}.$
+$\kappa_{n}(\mu_{p})=1\mbox{ if $n=0$ or $2$}\quad\quad\mathrm{and}\quad\quad\kappa_{n}(\mu_{p})=0\mbox{ otherwise}.$
 The number $F_{p}(k)=\frac{1}{pk+1}\binom{pk+1}{k}$ is the $k$-th Fuss-Catalan number of order $p$. They extend the Catalan numbers for $p\geq 2$.
 Theorem 3 (Free Poisson law in high order).
-For $p\geq 2$, the measure $\nu_{p,t}:=\mu_{b_{p,t}}$ is a compactly supported probability measure on $\mathbb{R}$ having for moments
-    if $p$ is even,
-    $m_{n}(\nu_{p,t})=\sum_{b=1}^{n}\frac{1}{b}\binom{n-1}{b-1}\binom{pn/2}{b-1}t^{b},$
-    if $p$ is odd,
-    $m_{n}(\nu_{p,t})=\sum_{b=1}^{n}\frac{1}{b}\binom{n-1}{b-1}\binom{pn/2}{b-1}t^{b}\mbox{ if $n$ even}\quad\quad\mathrm{and}\quad\quad m_{n}(\nu_{p,t})=0\mbox{if $n$ odd},$
+For $p\geq 1$, the measure $\nu_{2p,t}:=\mu_{b_{2p,t}}$ is a compactly supported probability measure on $\mathbb{R}$ having for moments
+$m_{n}(\nu_{2p,t})=\sum_{b=1}^{n}\frac{1}{b}\binom{n-1}{b-1}\binom{pn}{b-1}t^{b},$
 and for free cumulants
-    if $p$ is even,
-    $\kappa_{n}(\nu_{p,t})=t\mbox{ for all $n\geq 1$},$
-    if $p$ is odd,
-    $\kappa_{n}(\nu_{p,t})=t\mbox{ for all even $n\geq 1$}\quad\quad\mathrm{and}\quad\quad m_{n}(\nu_{p,t})=0\mbox{ for all odd $n\geq 1$}.$
+$\kappa_{n}(\nu_{2p,t})=t\mbox{ for all $n\geq 1$},$
 The polynomial $F_{p,k}(t)=\sum_{b=1}^{n}\frac{1}{b}\binom{k-1}{b-1}\binom{pk}{b-1}t^{b}$ is called the $k+1$-th Fuss-Narayana polynomial of order $p$ and parameter $t$.
@@
 $\mathbf{W}_{N}:=\frac{X}{N^{\frac{p-1}{2}}},$
-where $X$ is a real symmetric tensor with entries mean 00, variance $\frac{1}{(p-1)!}$ and bounded moments.
-This is in particular the case when the entries are Gaussian.
-On the other hand, a Wishart tensor is a real symmetric tensor of the form
-$\mathcal{W}^{k}_{N}:=\frac{x_{1}^{\underline{\otimes}2}+\ldots+x_{k}^{\underline{\otimes}2}}{N^{\frac{p}{2}}},$
-where the $x_{i}$ are $(p/2)$-order tensors with i.i.d entries with mean 00, variance $\frac{1}{[(p/2)!]^{2/p}}$ and bounded moments (it is for even $p$, see odd case and formal definitions in Section 3.1).
+where $X$ is a real symmetric $p$-order tensor having entries with mean 00, variance $\frac{1}{(p-1)!}$ and bounded moments.
+This is in particular the case when $X$ belongs to the Gaussian Orthogonal Tensor Ensemble.
+On the other hand, a $2p$-order Wishart tensor is a real symmetric tensor of the form
+$\mathcal{W}^{k}_{N}:=\frac{x_{1}^{\underline{\otimes}2}+\ldots+x_{k}^{\underline{\otimes}2}}{N^{p}},$
+where the $x_{i}$ are $p$-order tensors with i.i.d entries with mean 00, variance $\frac{1}{(p!)^{1/p}}$ and bounded moments and $x_{1}^{\underline{\otimes}2}$ is a symmetrization of $x_{i}\otimes x_{i}$.
+All of that is for even $p$, see the odd case and formal definitions in Section 3.1.
 We then have the following convergences.
@@
 $\mathcal{W}$ and $b$).
-The last convergence result we state is the free Central Limit Theorem for tensors.
+The last convergence result we state now is the free Central Limit Theorem for tensors.
 Theorem 6 (Free CLT).
 Let $p\geq 2$ be a fixed even integer and $(T_{i})_{i\geq 1}$ a collection of tensors of order $p$ such that
-$m_{1}(T)=0\quad\quad\mathrm{and}\quad\quad m_{2}(T)=1\quad\quad\mathrm{and}\quad\quad\forall i,\forall\mathfrak{b}\in\mathcal{M}_{0},|\mathfrak{b}(T_{i})|\leq C(\mathfrak{b}).$
+$m_{1}(T_{i})=0\quad\quad\mathrm{and}\quad\quad m_{2}(T_{i})=1\quad\quad\mathrm{and}\quad\quad\forall i,\forall\mathfrak{b}\in\cup_{n}\mathcal{B}_{n},|\mathfrak{b}(T_{i})|\leq C(\mathfrak{b}).$
 Denote $T^{k}=\frac{1}{\sqrt{k}}\sum_{i=1}^{k}T_{i}$. Then when $k$ goes to infinity,
@@
 $\nu_{p,t}\oplus_{p}\nu_{p,t^{\prime}}=\nu_{p,t+t^{\prime}}.$
-We can also define the $R$-transform which is additive for two freely independent tensors.
-Theorem 7.
+We may also define the $R$-transform which is additive for two freely independent tensors.
+Corollary 3.
 The $R$-transform and the $Q$-transform are defined as
@@
 It is where the $Q$-transform could be useful.
-These developments will be the subject of further works, jointly with a better understanding of the link between this measure and the $z$-eigenvalues of a given tensor.
+These developments will be the subject of further works, jointly with a better understanding of the link between this measure and the eigenvalues of a given tensor, which is still unclear.
 These results also open paths to treat some questions about high dimension statistics for random tensors.
@@
 In Section 2, we will introduce some formalism and prove the analytic moment-cumulant formula.
-In Section 3, we will define Wigner and Wishart and prove the convergence of their associated laws.
+In Section 3, we will define Wigner and Wishart tensors and prove the convergence of their associated laws.
 In Section 4, we will finally introduce tensorial free convolution for compactly supported measures, $R$-transform and give some basic examples.
@@
 For a given choice of inputs and outputs edges, there are $[(p/2)!]^{n}$ such maps.
-For $p$ odd, the definition of an odd multicycle, presented in the introduction follows.
+For $p$ odd, an odd multicycle is a map with $2n$ vertices and $(p+1)/2$ edges between the $12i-$ 1-th and the $2i$-th ones, $(p-1)/2$ edges between the $2i$-th and the $12i+$ 1-th ones, for all $1\leq i\leq n$ ($12n+1=$ 1).
 $M$$\frac{1}{N}\mathrm{Tr}(M)$
@@
 We denote $\mathfrak{b}(T)$ instead of $\mathfrak{b}(T,\cdots,T)$ for ease of notations, and the distribution of $T$ is then the collection of the $\mathfrak{b}(T)$ for all $p$-regular trace map.
-We define $\mathcal{B}_{n}$ the set of connected rooted $p$-regular combinatorial maps in $\mathcal{M}_{0}$ with $n$ (unlabeled) vertices.
+We remind thet $\mathcal{B}_{n}$ is the set of connected rooted $p$-regular combinatorial maps in $\mathcal{M}_{0}$ with $n$ (unlabeled) vertices.
 Definition 1 (Moments of a tensor).